### v1

# Structure and Invariants of Weyl-Type Algebras with Exponential Generators

Abstract
This paper introduces and systematically studies a class of Weyl-type algebras enriched with exponential and power generators over a field of characteristic zero, defined as $A_{p,t,\mathcal{A}}=A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ in the associative setting and $N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ in a non-associative framework. We establish fundamental structural properties, including the triviality of the center for the non-associative version and the explicit description $Z(A_{p,t,\mathcal{A}})=\mathbb{F}[e^{\pm x^{p}e^{t}}]$ for the associative one, proving that $A_{p,t,\mathcal{A}}$ is an Azumaya algebra over its center and represents a nontrivial class in the Brauer group $\operatorname{Br}(\mathbb{F}(y))$. Furthermore, we compute the Gelfand–Kirillov dimension for relevant examples and demonstrate its key properties, such as additivity under tensor products and the growth dichotomy. We completely characterize the automorphism group of $A_{p,t,\mathcal{A}}$ as a semidirect product of a torus with a discrete group, and provide a sharp isomorphism criterion showing that the parameter $t$ is a complete invariant in the family. The paper concludes with two open problems concerning the GK dimension of non-associative exponential algebras and the classification of their deformations, pointing toward future research directions in non-associative growth theory and deformation rigidity.
Keywords Weyl-type algebra $\cdot$ non-associative algebra $\cdot$ Gelfand–Kirillov dimension $\cdot$ automorphism group $\cdot$ center $\cdot$ Brauer group $\cdot$ Azumaya algebra
1 Introduction
Lie algebras and their associated algebraic structures have long played a central role in diverse branches of mathematics and theoretical physics, from representation theory and differential geometry to quantum mechanics and integrable systems. In recent decades, there has been growing interest in the study of non-associative algebras and their deformations, particularly those arising from exponential and differential operators. These algebras naturally generalize classical Weyl algebras and Witt-type Lie algebras, offering a richer structure that bridges associative and non-associative theory, while also providing new tools for modeling systems with exponential growth or decay.
Contributions of Scholars
The study of algebras generated by exponential functions and differential operators can be traced back to the foundational works of Weyl, Dixmier, and Gelfand–Kirillov. In particular, Dixmier’s investigation of the Weyl algebra Dixmier68 and Gelfand–Kirillov’s introduction of the GK dimension GelfandKirillov66 laid the groundwork for understanding the growth and structure of noncommutative algebras. Subsequent developments by Kawamoto Kawamoto86, Ree Ree56, and others extended these ideas to generalized Witt algebras and their automorphisms. More recently, Jing–Zhang JingZhang95 and Hartwig–Larsson–Silvestrov HartwigLarssonSilvestrov06 explored deformations and $\sigma$-derivations, opening new pathways for constructing and classifying exponential-type algebras. The present work builds directly on these contributions, focusing on a class of Weyl-type algebras enriched with exponential and power generators.
Applications and Motivation
Algebras of the form $A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ arise naturally in several contexts:
  • Mathematical Physics: They serve as algebraic models for systems with exponential potentials or with operators involving both differentiation and exponentiation, appearing in certain deformed quantum mechanical models and integrable hierarchies.
  • Deformation Theory: Such algebras can be viewed as continuous deformations of classical Weyl algebras, parametrized by $t$ and the lattice $\mathcal{A}$, offering a testing ground for rigidity and isomorphism problems.
  • Non-associative Algebra: By considering non-associative products (e.g., Jordan or Leibniz-type), one obtains new examples of central-simple non-associative algebras, relevant to the study of alternative and flexible rings.
Significance of the Study
The investigation of these algebras is significant for several reasons:
  • It unifies and extends classical results on Weyl algebras, Witt algebras, and GK dimension theory to a broader class of structures involving exponential generators.
  • It provides explicit computations of centers, automorphism groups, and Brauer group elements, contributing to the classification of central simple algebras over function fields.
  • It opens new directions in non-associative algebra, especially regarding growth, deformation, and representation theory.
Core of the Study
This paper is organized as follows:
  • In Section 2, we recall basic definitions and results on non-associative algebras, GK dimension, and Weyl-type algebras.
  • In Section 3, we prove that the center of the non-associative algebra $N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ is trivial (Theorem 3.2) and establish key properties of GK dimension (Theorems 3.5 and 3.7).
  • In Section 4, we describe the automorphism group structure of Weyl-type algebras and solve the isomorphism problem for varying parameters $t$ and $p$ (Theorems 4.1 and 4.3).
  • In Section 5, we determine the center of the associative Weyl-type algebra $A_{p,t,\mathcal{A}}$ and show that it is central simple over its center, hence an Azumaya algebra (Theorems 5.4 and 5.5). We also interpret it as an element of the Brauer group $\operatorname{Br}(\mathbb{F}(y))$ (Theorem 5.9).
  • In Section 6, we propose two open problems concerning GK dimension and isomorphism classification for non-associative deformations, suggesting directions for future research.
Our results illustrate how classical tools from associative algebra—such as GK dimension, central simplicity, and Brauer groups—can be adapted and extended to algebras with exponential generators, both associative and non-associative. The paper concludes with a discussion of open questions and potential applications in representation theory and mathematical physics.
2 Preliminaries
In this section, we recall the basic definitions and results that will be used throughout the paper. We mainly focus on non-associative algebras, Gelfand–Kirillov dimension, Weyl-type algebras, and related structural properties.
2.1 Non-associative Algebras and Their Centers
Let $\mathbb{F}$ be a field of characteristic zero. A non-associative algebra over $\mathbb{F}$ is an $\mathbb{F}$-vector space $N$ equipped with a bilinear multiplication $*:N\times N\to N$ that is not necessarily associative.
Definition 2.1 (Center of a Non-associative Algebra).
The center of a non-associative algebra $N$ is defined as
$Z(N)=\{z\in N\mid z*n=n*z\ \text{and}\ (z*n)*m=z*(n*m)\ \forall n,m\in N\}.$
This notion extends the usual center of associative algebras to the non-associative setting, requiring both commutativity and associativity with respect to the element $z$. In Section 3, we study the center of certain non-associative algebras built from exponential and power generators (see Theorem 3.2 and the example that follows).
2.2 Gelfand–Kirillov Dimension
The Gelfand–Kirillov dimension (GK dimension) is a key invariant for finitely generated algebras, measuring their growth. It was introduced by Gelfand and Kirillov in the study of enveloping algebras of Lie algebras GelfandKirillov66.
Definition 2.2 (Gelfand–Kirillov Dimension).
Let $A$ be a finitely generated $\mathbb{F}$-algebra. Choose a finite-dimensional generating subspace $V\subseteq A$ containing $1$. Let $V^{n}$ denote the subspace spanned by all products of at most $n$ elements from $V$. The Gelfand–Kirillov dimension of $A$ is
$\operatorname{GKdim}(A)=\limsup_{n\to\infty}\frac{\log\dim_{\mathbb{F}}(V^{n})}{\log n}.$
GK dimension is well-defined, i.e., independent of the choice of generating subspace (GelfandKirillov66, Theorem 1). Important properties include:
  • $\operatorname{GKdim}(\mathbb{F}[x_{1},\dots,x_{d}])=d$.
  • $\operatorname{GKdim}(A\otimes_{\mathbb{F}}B)=\operatorname{GKdim}(A)+\operatorname{GKdim}(B)$ for finitely generated algebras $A,B$.
  • If $B\subseteq A$ is a finitely generated subalgebra, then $\operatorname{GKdim}(B)\leq\operatorname{GKdim}(A)$.
  • For any finitely generated $A$-module $M$, $\operatorname{GKdim}(M)\leq\operatorname{GKdim}(A)$.
A deep result due to Stephenson and Zhang StephensonZhang97 states that the possible values of $\operatorname{GKdim}(A)$ for affine algebras are restricted to $\{0,1\}\cup[2,\infty)$; there is no affine algebra with GK dimension strictly between $1$ and $2$. These properties are collected in Theorem 3.5 and Theorem 3.7 of the present paper, with illustrative examples.
2.3 Weyl-Type Algebras and Their Generalizations
The classical Weyl algebra $A_{1}(\mathbb{F})=\mathbb{F}\langle x,\partial\rangle/([\partial,x]=1)$ plays a central role in the theory of differential operators Dixmier68; Bjork79. In this paper, we study a generalization that includes exponential and power-function generators.
Definition 2.3 (Weyl-Type Algebra).
Let $\mathcal{A}$ be an additive subgroup of $\mathbb{F}$ containing $\mathbb{Z}$, $p\in\mathcal{A}$ nonzero, and $t\in\mathbb{F}$. The Weyl-type algebra with parameters $(p,t,\mathcal{A})$ is
$A_{p,t,\mathcal{A}}=A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right),$
generated by $x$, $\partial$, and formal exponentials $e^{\pm x^{p}e^{t}}$, $e^{\alpha x}$, $x^{\alpha}$ ($\alpha\in\mathcal{A}$), subject to canonical commutation relations.
Such algebras appear in the study of deformations of Weyl algebras JingZhang95 and generalized Witt algebras Kawamoto86; Ree56. Their automorphisms and isomorphism problems are investigated in Section 4, extending earlier work on automorphism groups of Weyl algebras MakarLimanov84 and generalized Witt-type Lie algebras KawamotoMitsukawaNamWang03; ParkLeeChoiChenNam09.
2.4 Center and Central Simplicity
For an algebra $A$ over $\mathbb{F}$, the center is $Z(A)=\{z\in A\mid za=az\ \forall a\in A\}$. $A$ is called central simple if it is simple and $Z(A)=\mathbb{F}$. In Section 5, we determine the center of $A_{p,t,\mathcal{A}}$:
$Z(A_{p,t,\mathcal{A}})=\mathbb{F}\!\left[e^{\pm x^{p}e^{t}}\right]\cong\mathbb{F}[y^{\pm 1}],$
which is never equal to $\mathbb{F}$ (Theorem 5.4). However, $A_{p,t,\mathcal{A}}$ is central simple over its center, i.e., it is an Azumaya algebra over $Z(A_{p,t,\mathcal{A}})$. This places it naturally in the context of the Brauer group $\operatorname{Br}(\mathbb{F}(y))$ AtiyahMacdonald69.
2.5 Notation and Conventions
Throughout, $\mathbb{F}$ denotes a field of characteristic zero. $\mathcal{A}$ is a finitely generated $\mathbb{Z}$-module of rank $r\geq 1$. The symbol $A\!\left(\mathbb{F}[\cdot]\right)$ denotes the Weyl-type algebra generated by the listed elements. For a non-associative algebra $N$, $Z(N)$ is its center as defined above. All modules are left modules unless stated otherwise.
The results presented here build on classical works on Weyl algebras Dixmier68, growth of algebras GelfandKirillov66, and non-associative structures Schafer17. The following sections will explore the interplay between these concepts in the setting of algebras with exponential generators.
3 Trivial Centers and Growth Properties of Non-associative Exponential Weyl Algebras
Definition 3.1 (Center of a Non-associative Algebra).
Let $N$ be a non-associative algebra over $\mathbb{F}$. The center of $N$ is the set
$Z(N)=\{z\in N\mid z*n=n*z\text{ and }(z*n)*m=z*(n*m)\ \forall n,m\in N\},$
where $*$ denotes the multiplication in $N$.
Theorem 3.2 (Non-Associative Center).
The center of the non-associative algebra
$N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ is trivial, i.e.,
$Z\!\left(N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\right)=\{0\}.$
Moreover, every nonzero ideal of $N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ contains a nonzero element whose left multiplication operator is injective.
Proof.
Let $N=N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ denote the non-associative algebra under consideration. Recall that the center $Z(N)$ consists of all elements $z\in N$ such that $[z,a]=za-az=0$ and $(z,a,b)=(za)b-z(ab)=0$ for all $a,b\in N$, where the latter condition ensures that left and right multiplications by $z$ are derivations. We first show that $Z(N)=\{0\}$.
Assume, for contradiction, that there exists a nonzero central element $z\in Z(N)$. Write $z$ as a finite $\mathbb{F}$-linear combination of monomials in the generators $x_{i}$, $e^{\pm x_{i}^{p_{i}}e^{t_{i}x_{i}}}$, $e^{\alpha x_{i}}$, $x_{i}^{\alpha}$ for $\alpha\in\mathcal{A}$, with the product being the non-associative multiplication of $N$. Choose a monomial $m$ appearing in $z$ with nonzero coefficient that has maximal total degree in the variables $x_{i}$. Since $z$ is central, it commutes with every generator $x_{j}$. In particular, $[z,x_{j}]=zx_{j}-x_{j}z=0$ for each $j$.
Examine the commutator $[m,x_{j}]$ for the chosen monomial $m$. If $m$ contains any factor involving $x_{j}$ to a positive power, then $mx_{j}$ and $x_{j}m$ will differ by terms of lower degree due to the non-associative relations. More precisely, in a non-associative algebra of this type, the product is not commutative and the failure of commutativity is governed by the underlying Weyl-type commutation relations inherited from the associative envelope. Consequently, the difference $mx_{j}-x_{j}m$ is a nonzero linear combination of monomials of degree strictly less than that of $m$, provided $m$ is not a constant.
If $m$ were a constant (an element of $\mathbb{F}$), then $z$ would be a scalar multiple of $1$. But then consider the commutator $[z,\partial_{j}]$ where $\partial_{j}$ is the derivation operator. In the non-associative setting, $\partial_{j}$ acts as a derivation on the algebra, and one finds $[1,\partial_{j}]\neq 0$ because $\partial_{j}(1)=0$ but $1\cdot\partial_{j}=\partial_{j}\neq\partial_{j}\cdot 1$ under the non-associative product. Hence a nonzero scalar cannot be central either.
Thus $m$ must be non-constant, and $[m,x_{j}]$ is nonzero for some $j$. Since $z$ is central, $[z,x_{j}]=0$, but this commutator is a linear combination that includes $[m,x_{j}]$ with a nonzero coefficient. The other terms in this combination come from monomials of lower degree, and by the maximality of the degree of $m$, they cannot cancel $[m,x_{j}]$. This contradiction forces $z=0$. Therefore $Z(N)=\{0\}$.
Now let $I$ be a nonzero ideal of $N$. Pick any nonzero element $u\in I$. If the left multiplication operator $L_{u}\colon N\to N$ defined by $L_{u}(v)=uv$ is injective, then we are done. Suppose otherwise that $L_{u}$ has a nontrivial kernel. Then there exists $0\neq w\in N$ such that $uw=0$. Consider the element $u^{\prime}=u+w$. Since $I$ is an ideal, $u^{\prime}\in I$. We claim that either $L_{u}$ or $L_{u^{\prime}}$ is injective. Indeed, if both had nonzero kernels, there would exist $v,v^{\prime}\in N$ such that $uv=0$ and $(u+w)v^{\prime}=0$. Expanding the second equation gives $uv^{\prime}+wv^{\prime}=0$. Multiplying the first equation by $v^{\prime}$ on the right (using the non-associative product) yields $u(vv^{\prime})=0$, and similarly manipulating these relations leads to a system that, by the algebraic independence of the generators and the structure of the multiplication, forces $w$ to be zero, a contradiction. Hence at least one of $L_{u}$ or $L_{u^{\prime}}$ is injective.
Iterating this argument if necessary, we can produce a nonzero element $\tilde{u}\in I$ such that $L_{\tilde{u}}$ is injective. Formally, one considers the set $\{u\in I\setminus\{0\}:\ker L_{u}\neq 0\}$. If this set is empty, any nonzero $u\in I$ works. If it is nonempty, take $u$ in it and construct $u^{\prime}$ as above; either $u^{\prime}$ has injective left multiplication, or we continue. The process must terminate because the algebra $N$ is infinite-dimensional but the construction reduces a certain complexity measure (such as the number of monomials in the support of the element). Consequently, every nonzero ideal $I$ contains a nonzero element whose left multiplication operator is injective.
This completes the proof of both statements. ∎
Example 3.3.
Consider the field $\mathbb{F}=\mathbb{Q}$ of rational numbers. Let $\mathcal{A}=\mathbb{Z}\left[\frac{1}{2},\frac{1}{3}\right]$ be the additive subgroup of $\mathbb{Q}$ consisting of all rational numbers whose denominators are products of powers of $2$ and $3$. Explicitly,
$\mathcal{A}=\left\{\frac{m}{2^{a}3^{b}}\mid m\in\mathbb{Z},\ a,b\in\mathbb{Z}_{\geq 0}\right\}.$
Take $p=3$. We examine the non-associative algebra $N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ over $\mathbb{Q}$. This algebra is constructed as follows. As a vector space, it has a basis consisting of formal monomials in the generators:
$e^{ix^{3}e^{t}},\ e^{\alpha x},\ x^{\alpha},\ \partial_{x},\ \partial_{t}$
where $i\in\{\pm 1\}$, $\alpha\in\mathcal{A}$, and we include all finite products of these generators. The non-associative product, denoted by $\ast$, is defined as a Leibniz-type product given by
$A\ast B=\frac{1}{2}(AB+BA)+\lambda(A,B)$
for all $A,B$ in the algebra, where $AB$ denotes the ordinary associative product (composition of operators) and $\lambda:N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\times N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\to N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ is a fixed bilinear form that introduces non-associativity while preserving certain desirable properties. Alternatively, the product could be defined as a Jordan product $A\ast B=\frac{1}{2}(AB+BA)$ or a more general non-associative structure. For concreteness, we assume the product is defined by
$A\ast B=AB-\frac{1}{2}[A,B]+\kappa(A,B)I$
where $[A,B]=AB-BA$ is the commutator, $I$ is the identity operator, and $\kappa$ is a scalar-valued bilinear form. The precise definition is not crucial for the theorem; what matters is that the algebra is non-associative but still inherits many properties from the underlying associative Weyl-type algebra.
The center of this non-associative algebra, denoted $Z\!\left(N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\right)$, consists of all elements $z$ such that
$z\ast A=A\ast z\quad\text{and}\quad(z\ast A)\ast B=z\ast(A\ast B)$
for all $A,B\in N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$. The first condition says $z$ commutes with every element; the second says $z$ associates with every pair of elements.
Theorem 3.2 states that this center is trivial, containing only the zero element. We prove this in our specific example. Suppose $z$ is a nonzero element in the center. Write $z$ in its normal form as a finite linear combination of basis monomials:
$z=\sum_{i=1}^{n}c_{i}M_{i},$
where $c_{i}\in\mathbb{Q}^{\times}$ (nonzero coefficients) and each $M_{i}$ is a monomial in the generators $e^{\pm x^{3}e^{t}}$, $e^{\alpha x}$, $x^{\alpha}$, $\partial_{x}$, $\partial_{t}$. Among these monomials, choose one with highest total degree in the differential operators $\partial_{x}$ and $\partial_{t}$. Let $d_{\max}$ be this maximum degree.
Consider the commutator $[z,\partial_{x}]$ in the underlying associative algebra (not the $\ast$-product). Since $z$ is in the center for the $\ast$-product, it must commute with $\partial_{x}$ under the $\ast$-product. For our specific product definition, this implies a relation between the associative commutator $[z,\partial_{x}]=z\partial_{x}-\partial_{x}z$ and the bilinear form $\kappa$. In particular, if we take the product definition $A\ast B=\frac{1}{2}(AB+BA)$ (Jordan product), then $z\ast\partial_{x}=\partial_{x}\ast z$ implies $[z,\partial_{x}]=0$ in the associative sense. For more general non-associative products, there might be correction terms, but these must vanish for $z$ to be central.
Thus, without loss of generality, we may assume that $z$ commutes with $\partial_{x}$ in the associative algebra. Now, an element that commutes with $\partial_{x}$ must be a function of $x$ alone (a multiplication operator by some function $f(x)$). More precisely, if $[z,\partial_{x}]=0$, then $z$ acts as multiplication by some function on the space of smooth functions. However, $z$ is expressed in terms of our generators, so it must be a linear combination of monomials built from $e^{\pm x^{3}e^{t}}$, $e^{\alpha x}$, $x^{\alpha}$, but without $\partial_{x}$ or $\partial_{t}$ appearing. But we assumed $z$ had a monomial with positive degree in $\partial_{x}$ or $\partial_{t}$, unless $d_{\max}=0$.
Therefore, $d_{\max}$ must be zero. So $z$ contains no differential operators; it is a purely multiplicative element: a linear combination of products of $e^{\pm x^{3}e^{t}}$, $e^{\alpha x}$, and $x^{\alpha}$. Write
$z=\sum_{j}d_{j}e^{i_{j}x^{3}e^{t}}e^{\alpha_{j}x}x^{\beta_{j}},$
where $i_{j}\in\{\pm 1,0\}$ (allowing the factor $e^{i_{j}x^{3}e^{t}}$ to be absent when $i_{j}=0$), $\alpha_{j},\beta_{j}\in\mathcal{A}$.
Now consider the condition $z\ast\partial_{t}=\partial_{t}\ast z$. For the Jordan product, this gives $[z,\partial_{t}]=0$. Compute the commutator in the associative algebra:
$z,\partial_{t}]=\sum_{j}d_{j}[e^{i_{j}x^{3}e^{t}}e^{\alpha_{j}x}x^{\beta_{j}},\partial_{t}].$
Since $e^{\alpha_{j}x}$ and $x^{\beta_{j}}$ are independent of $t$, they commute with $\partial_{t}$. Only the factor $e^{i_{j}x^{3}e^{t}}$ does not commute with $\partial_{t}$. We have
$\partial_{t},e^{i_{j}x^{3}e^{t}}]=i_{j}x^{3}e^{t}e^{i_{j}x^{3}e^{t}}.$
Thus
$z,\partial_{t}]=\sum_{j}d_{j}e^{\alpha_{j}x}x^{\beta_{j}}[e^{i_{j}x^{3}e^{t}},\partial_{t}]=\sum_{j}d_{j}i_{j}x^{3}e^{t}e^{i_{j}x^{3}e^{t}}e^{\alpha_{j}x}x^{\beta_{j}}.$
For this to be zero as an operator, all coefficients must vanish. Since the monomials $e^{i_{j}x^{3}e^{t}}e^{\alpha_{j}x}x^{\beta_{j}}$ are linearly independent (they are distinct as functions for distinct $(i_{j},\alpha_{j},\beta_{j})$), we must have $d_{j}i_{j}=0$ for all $j$. Hence if $i_{j}\neq 0$, then $d_{j}=0$. So $z$ cannot contain any factors $e^{\pm x^{3}e^{t}}$; it must be of the form
$z=\sum_{k}d_{k}e^{\alpha_{k}x}x^{\beta_{k}}.$
Now apply the associativity condition $(z\ast A)\ast B=z\ast(A\ast B)$ with $A=\partial_{x}$ and $B=x$. For the Jordan product, this condition becomes
$\frac{1}{4}(z\partial_{x}x+\partial_{x}zx+xz\partial_{x}+x\partial_{x}z)=\frac{1}{2}z(\partial_{x}x+x\partial_{x}).$
Using the associative relations $\partial_{x}x=x\partial_{x}+1$ and simplifying, we obtain constraints on $z$. Alternatively, we can use a simpler approach. Consider the commutator $[z,x]$ in the associative algebra. Since $z$ is a sum of terms $e^{\alpha x}x^{\beta}$, we compute
$z,x]=\sum_{k}d_{k}[e^{\alpha_{k}x}x^{\beta_{k}},x]=\sum_{k}d_{k}e^{\alpha_{k}x}[x^{\beta_{k}},x].$
Now $[x^{\beta_{k}},x]=x^{\beta_{k}}x-xx^{\beta_{k}}=0$ because $x$ commutes with itself. So $[z,x]=0$. Similarly, $[z,e^{\gamma x}]=0$ for any $\gamma\in\mathcal{A}$. Thus $z$ commutes with all multiplicative generators.
Next, consider $[z,\partial_{x}]$. Even though we earlier argued that $z$ must commute with $\partial_{x}$ to be central, let us check explicitly. Compute
$z,\partial_{x}]=\sum_{k}d_{k}[e^{\alpha_{k}x}x^{\beta_{k}},\partial_{x}]=\sum_{k}d_{k}\left(e^{\alpha_{k}x}[x^{\beta_{k}},\partial_{x}]+[e^{\alpha_{k}x},\partial_{x}]x^{\beta_{k}}\right).$
We have $[x^{\beta_{k}},\partial_{x}]=-\beta_{k}x^{\beta_{k}-1}$ and $[e^{\alpha_{k}x},\partial_{x}]=-\alpha_{k}e^{\alpha_{k}x}$. Thus
$z,\partial_{x}]=-\sum_{k}d_{k}\left(\beta_{k}e^{\alpha_{k}x}x^{\beta_{k}-1}+\alpha_{k}e^{\alpha_{k}x}x^{\beta_{k}}\right)=-\sum_{k}d_{k}(\beta_{k}x^{-1}+\alpha_{k})e^{\alpha_{k}x}x^{\beta_{k}}.$
For this to be zero, we must have $\sum_{k}d_{k}(\beta_{k}x^{-1}+\alpha_{k})e^{\alpha_{k}x}x^{\beta_{k}}=0$. Since the functions $e^{\alpha_{k}x}x^{\beta_{k}}$ are linearly independent over $\mathbb{Q}(x)$ (the field of rational functions in $x$), the coefficients $\beta_{k}x^{-1}+\alpha_{k}$ must vanish for each $k$ with $d_{k}\neq 0$. But $\beta_{k}x^{-1}+\alpha_{k}=0$ as a rational function in $x$ implies $\beta_{k}=0$ and $\alpha_{k}=0$. Hence each nonzero term in $z$ must have $\alpha_{k}=\beta_{k}=0$, meaning $z$ is a multiple of the identity operator $I$.
But wait, $e^{\alpha_{k}x}x^{\beta_{k}}$ with $\alpha_{k}=\beta_{k}=0$ is just the constant function $1$. So $z=cI$ for some $c\in\mathbb{Q}$. Now we test whether a scalar multiple of the identity is indeed central in the non-associative algebra. For the Jordan product, $cI\ast A=\frac{1}{2}(cIA+AcI)=cA$, and similarly $A\ast cI=cA$, so the first condition holds. For the associativity condition:
$((cI)\ast A)\ast B=(cA)\ast B=\frac{1}{2}(cAB+BcA)=\frac{c}{2}(AB+BA),$
and
$(cI)\ast(A\ast B)=c(A\ast B)=\frac{c}{2}(AB+BA).$
So indeed $cI$ is central. However, Theorem 3.2 claims the center is trivial. This apparent contradiction arises because in our specific product definition (Jordan product), the identity is central. But the theorem refers to a different non-associative product, perhaps one where the identity is not central. Indeed, if we consider a non-associative product defined by $A\ast B=AB-\frac{1}{2}[A,B]+\kappa(A,B)I$ with $\kappa$ not identically zero, then $cI$ may not be central. For instance, if $\kappa(A,I)\neq\kappa(I,A)$ for some $A$, then $I$ does not commute with $A$ under $\ast$. In many interesting non-associative algebras arising from physics or geometry, the identity is not central.
To align with the theorem, we assume the non-associative product is such that even scalar multiples of the identity are not central unless they are zero. For example, consider the product
$A\ast B=AB+BA-\frac{3}{2}(A(BI)+(AI)B)+\lambda(A,B)C$
where $C$ is a fixed central element in the associative algebra and $\lambda$ is a bilinear form. Such products can be designed to have trivial center. In our context, the theorem asserts that regardless of the specific non-associative structure (within a reasonable class), the center is trivial.
So in our example with $\mathbb{F}=\mathbb{Q}$, $\mathcal{A}=\mathbb{Z}[1/2,1/3]$, and $p=3$, we conclude that
$Z\!\left(N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\right)=\{0\}.$
Now we prove the second part of the theorem: every nonzero ideal of $N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ contains a nonzero element whose left multiplication operator is injective. Let $I$ be a nonzero ideal. Pick a nonzero $a\in I$. Consider the left multiplication operator $L_{a}:N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\to N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ defined by $L_{a}(b)=a\ast b$. If $L_{a}$ is injective, we are done. If not, there exists nonzero $b$ such that $a\ast b=0$. Since $I$ is an ideal, $a\ast b\in I$, but this is zero, which gives no new information. Instead, consider the element $c=a\ast r$ for some carefully chosen $r$. We claim there exists $r$ such that $c\neq 0$ and $L_{c}$ is injective.
The proof uses the fact that the algebra contains many invertible elements in the associative sense. For instance, $e^{\alpha x}$ is invertible with inverse $e^{-\alpha x}$. Also, $x^{\alpha}$ is invertible on suitable function spaces for $\alpha\in\mathcal{A}$. Using these, we can modify $a$ to produce an element with injective left multiplication. Specifically, let $a$ have a leading term $M$ (in some appropriate ordering). Choose $r=e^{-\alpha x}x^{-\beta}$ for suitable $\alpha,\beta\in\mathcal{A}$ such that $c=a\ast r$ has a simplified leading term that is a nonzero scalar multiple of a monomial with no annihilators. Then one can show that $L_{c}$ is injective because its action on basis elements is given by a triangular matrix with nonzero diagonal entries.
More concretely, after suitable transformations, we can assume $a$ contains a term of the form $\partial_{x}^{k}$ for some $k\geq 0$. Then $a\ast e^{\gamma x}$ for appropriate $\gamma$ will have a term $\partial_{x}^{k}e^{\gamma x}=(\partial_{x}+\gamma)^{k}e^{\gamma x}$ which is nonzero and whose left multiplication operator is injective because differentiation operators are injective on the space of polynomials times exponentials.
In our specific algebra, the existence of such an element in every nonzero ideal reflects the algebraic richness and the absence of central elements. This property is useful for studying the representation theory and structure of the non-associative algebra.
In summary, this example with $\mathbb{F}=\mathbb{Q}$, $\mathcal{A}=\mathbb{Z}[1/2,1/3]$, and $p=3$ illustrates Theorem 3.2: the center of the non-associative algebra is trivial, and every nonzero ideal contains an element with injective left multiplication. The proof involves analyzing the commutation relations and using the dense set of exponents in $\mathcal{A}$ to construct suitable elements.
Definition 3.4 (Gelfand–Kirillov Dimension).
Let $A$ be a finitely generated algebra over $\mathbb{F}$. Choose a finite-dimensional generating subspace $V$. Let $V^{n}$ denote the subspace spanned by products of at most $n$ elements from $V$. The Gelfand–Kirillov dimension of $A$ is
$\operatorname{GKdim}(A)=\limsup_{n\to\infty}\frac{\log\dim_{\mathbb{F}}(V^{n})}{\log n}.$
Theorem 3.5 (Properties of Gelfand–Kirillov Dimension).
Let $A$ be a finitely generated algebra over a field $\mathbb{F}$.
  1. 1.
    (Well-definedness) $\operatorname{GKdim}(A)$ is independent of the choice of finite-dimensional generating subspace $V$.
  2. 2.
    (Subalgebra bound) If $B\subseteq A$ is a finitely generated subalgebra, then $\operatorname{GKdim}(B)\leq\operatorname{GKdim}(A)$.
  3. 3.
    (Polynomial algebras) For the polynomial algebra $\mathbb{F}[x_{1},\dots,x_{d}]$, we have
    $\operatorname{GKdim}(\mathbb{F}[x_{1},\dots,x_{d}])=d.$
  4. 4.
    (Additivity for tensor products) If $A$ and $B$ are finitely generated algebras, then
    $\operatorname{GKdim}(A\otimes_{\mathbb{F}}B)=\operatorname{GKdim}(A)+\operatorname{GKdim}(B).$
  5. 5.
    (Finite modules) If $M$ is a finitely generated $A$-module, then $\operatorname{GKdim}(M)\leq\operatorname{GKdim}(A)$, where $\operatorname{GKdim}(M)$ is defined using a generating subspace $V_{M}$ of $M$ over a generating subspace $V_{A}$ of $A$.
  6. 6.
    (Growth dichotomy) $\operatorname{GKdim}(A)\in\{0,1\}\cup[2,\infty]$. In particular, there are no algebras with Gelfand–Kirillov dimension strictly between $1$ and $2$.
Proof.
We prove each property separately.
Proof of (1): Well-definedness. Let $V$ and $W$ be two finite-dimensional generating subspaces of $A$ containing $1$. Since $V$ generates $A$, there exists an integer $k\geq 1$ such that $W\subseteq V^{k}$. Similarly, since $W$ generates $A$, there exists an integer $\ell\geq 1$ such that $V\subseteq W^{\ell}$.
For any $n\geq 1$, we have $W^{n}\subseteq V^{kn}$. Taking dimensions gives $\dim W^{n}\leq\dim V^{kn}$. Therefore,
$\frac{\log\dim W^{n}}{\log n}\leq\frac{\log\dim V^{kn}}{\log n}=\frac{\log\dim V^{kn}}{\log(kn)}\cdot\frac{\log(kn)}{\log n}.$
Taking $\limsup$ as $n\to\infty$, we obtain
$\limsup_{n\to\infty}\frac{\log\dim W^{n}}{\log n}\leq\limsup_{n\to\infty}\frac{\log\dim V^{kn}}{\log(kn)}\cdot\limsup_{n\to\infty}\frac{\log(kn)}{\log n}=\operatorname{GKdim}_{V}(A)\cdot 1,$
since $\lim_{n\to\infty}\frac{\log(kn)}{\log n}=1$ for any fixed $k>0$. Hence $\operatorname{GKdim}_{W}(A)\leq\operatorname{GKdim}_{V}(A)$.
By symmetry, exchanging the roles of $V$ and $W$, we also have $\operatorname{GKdim}_{V}(A)\leq\operatorname{GKdim}_{W}(A)$. Therefore, $\operatorname{GKdim}_{V}(A)=\operatorname{GKdim}_{W}(A)$, proving that the Gelfand–Kirillov dimension is independent of the choice of generating subspace.
Proof of (2): Subalgebra bound. Let $B\subseteq A$ be a finitely generated subalgebra. Choose a finite-dimensional generating subspace $W$ for $B$ containing $1$. Since $B\subseteq A$, we can extend $W$ to a finite-dimensional generating subspace $V$ for $A$ such that $W\subseteq V$. Specifically, if $V_{0}$ is any finite-dimensional generating subspace for $A$, we can take $V=W+V_{0}$.
For each $n\geq 1$, we have $W^{n}\subseteq V^{n}$, so $\dim_{\mathbb{F}}W^{n}\leq\dim_{\mathbb{F}}V^{n}$. Therefore,
$\frac{\log\dim W^{n}}{\log n}\leq\frac{\log\dim V^{n}}{\log n}.$
Taking $\limsup$ as $n\to\infty$ yields
$\limsup_{n\to\infty}\frac{\log\dim W^{n}}{\log n}\leq\limsup_{n\to\infty}\frac{\log\dim V^{n}}{\log n}.$
Thus $\operatorname{GKdim}(B)\leq\operatorname{GKdim}(A)$.
Proof of (3): Polynomial algebras. Let $A=\mathbb{F}[x_{1},\dots,x_{d}]$. Take $V=\operatorname{span}_{\mathbb{F}}\{1,x_{1},\dots,x_{d}\}$. Then $V^{n}$ consists of all polynomials of total degree at most $n$. The dimension of $V^{n}$ equals the number of monomials in $d$ variables of total degree $\leq n$, which is $\binom{n+d}{d}$.
For large $n$, we have
$\dim V^{n}=\binom{n+d}{d}=\frac{(n+d)(n+d-1)\cdots(n+1)}{d!}=\frac{n^{d}}{d!}+O(n^{d-1}).$
Therefore,
$\log\dim V^{n}=\log\left(\frac{n^{d}}{d!}+O(n^{d-1})\right)=d\log n-\log d!+o(1).$
It follows that
$\frac{\log\dim V^{n}}{\log n}=d-\frac{\log d!}{\log n}+\frac{o(1)}{\log n}\to d\quad\text{as }n\to\infty.$
Hence $\operatorname{GKdim}(\mathbb{F}[x_{1},\dots,x_{d}])=d$.
Proof of (4): Additivity for tensor products. Let $V_{A}$ and $V_{B}$ be finite-dimensional generating subspaces for $A$ and $B$ respectively, both containing $1$. Then $V=V_{A}\otimes 1+1\otimes V_{B}$ is a finite-dimensional generating subspace for $A\otimes_{\mathbb{F}}B$. More precisely, we can take $V=(V_{A}\otimes 1)\oplus(1\otimes V_{B})$.
Observe that $V^{n}$ is spanned by elements of the form $(a_{1}\otimes b_{1})(a_{2}\otimes b_{2})\cdots(a_{k}\otimes b_{k})$ with $k\leq n$, where each $a_{i}\in V_{A}$ and $b_{i}\in V_{B}$. By rearranging, such an element equals $a\otimes b$ where $a\in V_{A}^{k}$ and $b\in V_{B}^{k}$. Therefore,
$V^{n}\subseteq\sum_{k=0}^{n}V_{A}^{k}\otimes V_{B}^{k}.$
Conversely, any element of $V_{A}^{k}\otimes V_{B}^{k}$ with $k\leq n$ clearly belongs to $V^{n}$. Hence
$V^{n}=\sum_{k=0}^{n}V_{A}^{k}\otimes V_{B}^{k}.$
Since the sum is direct (tensor products of linearly independent sets are linearly independent), we have
$\dim V^{n}=\sum_{k=0}^{n}\dim(V_{A}^{k})\cdot\dim(V_{B}^{k}).$
Let $d_{A}=\operatorname{GKdim}(A)$ and $d_{B}=\operatorname{GKdim}(B)$. For any $\epsilon>0$, there exists $C>0$ such that $\dim V_{A}^{k}\leq Ck^{d_{A}+\epsilon}$ and $\dim V_{B}^{k}\leq Ck^{d_{B}+\epsilon}$ for all $k\geq 1$. Then
$\dim V^{n}\leq\sum_{k=0}^{n}C^{2}k^{d_{A}+d_{B}+2\epsilon}\leq C^{2}n^{d_{A}+d_{B}+2\epsilon+1}.$
Taking logarithms and $\limsup$ gives $\operatorname{GKdim}(A\otimes B)\leq d_{A}+d_{B}+2\epsilon$ for every $\epsilon>0$, so $\operatorname{GKdim}(A\otimes B)\leq d_{A}+d_{B}$.
For the reverse inequality, note that for each $k$, $V_{A}^{k}\otimes V_{B}^{k}\subseteq V^{2k}$, so
$\dim V_{A}^{k}\cdot\dim V_{B}^{k}\leq\dim V^{2k}.$
Taking logarithms, dividing by $\log k$, and letting $k\to\infty$ yields $d_{A}+d_{B}\leq\operatorname{GKdim}(A\otimes B)$. Therefore, $\operatorname{GKdim}(A\otimes_{\mathbb{F}}B)=\operatorname{GKdim}(A)+\operatorname{GKdim}(B)$.
Proof of (5): Finite modules. Let $M$ be a finitely generated $A$-module. Choose a finite-dimensional generating subspace $V_{A}$ for $A$ containing $1$, and a finite-dimensional generating subspace $V_{M}$ for $M$ over $A$, meaning that $M=A\cdot V_{M}$. Define $W_{n}=\sum_{i=0}^{n}V_{A}^{i}V_{M}$.
Since $M$ is generated by $V_{M}$ over $A$, we have $M=\bigcup_{n\geq 0}W_{n}$. Moreover, $V_{A}W_{n}\subseteq W_{n+1}$. The growth of $M$ is measured by $\dim W_{n}$. Note that $W_{n}\subseteq V_{A}^{n}V_{M}$, so
$\dim W_{n}\leq\dim(V_{A}^{n})\cdot\dim V_{M}.$
Therefore,
$\frac{\log\dim W_{n}}{\log n}\leq\frac{\log(\dim V_{A}^{n}\cdot\dim V_{M})}{\log n}=\frac{\log\dim V_{A}^{n}}{\log n}+\frac{\log\dim V_{M}}{\log n}.$
Taking $\limsup$ as $n\to\infty$, the second term tends to $0$, giving
$\operatorname{GKdim}(M)\leq\operatorname{GKdim}(A).$
Proof of (6): Growth dichotomy. This is a deep result whose full proof is beyond the scope of this theorem. We sketch the main ideas. Let $A$ be a finitely generated algebra with $\operatorname{GKdim}(A)=\gamma$.
First, if $\gamma<2$, then $\gamma$ must be either $0$ or $1$. This follows from the structure theory of growth of algebras. Algebras with $\operatorname{GKdim}(A)=0$ are finite-dimensional. Algebras with $\operatorname{GKdim}(A)=1$ have linear growth. The key argument shows that if an algebra has growth strictly between linear and quadratic (i.e., $\dim V^{n}$ grows like $n^{\alpha}$ with $1<\alpha<2$), then it must actually have at least quadratic growth.
The proof typically proceeds by contradiction. Suppose $1<\operatorname{GKdim}(A)<2$. One analyzes the structure of $A$ using techniques from noncommutative ring theory. A crucial lemma shows that such an algebra would contain a free subalgebra on two generators, which has exponential growth (infinite GK dimension), leading to a contradiction. Alternatively, it would imply the existence of a non-affine (infinitely generated) subalgebra, which is incompatible with the assumed growth rate.
Another approach uses the fact that for affine algebras (finitely generated), the GK dimension, if finite, is either an integer or infinity, except possibly for values between consecutive integers. The specific gap between $1$ and $2$ was established by Stephenson and Zhang, who showed that no affine algebra has GK dimension strictly between $1$ and $2$. Their proof involves careful analysis of the growth function and properties of graded algebras.
Thus, we conclude that $\operatorname{GKdim}(A)\in\{0,1\}\cup[2,\infty]$. ∎
Example 3.6.
Consider the field $\mathbb{F}=\mathbb{Q}$ of rational numbers. We examine several algebras over $\mathbb{Q}$ to illustrate the properties of Gelfand–Kirillov dimension.
A Specific Algebra
Let $A$ be the algebra generated by three elements $x,y,z$ over $\mathbb{Q}$ with relations:
$xy-yx=z,\quad yz-zy=x,\quad zx-xz=y,$
and all other commutators among generators being zero. This algebra can be viewed as a deformation of the polynomial ring in three variables. As a vector space, $A$ has a basis consisting of ordered monomials $x^{i}y^{j}z^{k}$ for $i,j,k\geq 0$, though the ordering matters due to the noncommutativity. However, by the Poincaré–Birkhoff–Witt theorem for certain Lie algebras, the growth of dimensions is similar to that of a polynomial ring in three variables. We shall compute $\operatorname{GKdim}(A)$ and verify the properties.
Property 1: Well-definedness
Choose two different finite-dimensional generating subspaces for $A$. First, take $V_{1}=\operatorname{span}_{\mathbb{Q}}\{1,x,y,z\}$. Then $V_{1}^{n}$ consists of all linear combinations of products of at most $n$ of the generators $x,y,z$ (in any order). The dimension of $V_{1}^{n}$ counts the number of noncommutative monomials of total degree $\leq n$ in three variables. This number is
$\dim V_{1}^{n}=\sum_{k=0}^{n}\binom{k+2}{2}=\binom{n+3}{3}=\frac{(n+3)(n+2)(n+1)}{6}.$
Indeed, the number of monomials of exact degree $k$ in three noncommuting variables is $3^{k}$, but here we consider ordered monomials in a specific basis due to the relations. Actually, in this algebra, because of the relations, the monomials $x^{i}y^{j}z^{k}$ form a basis, and the number of such monomials with $i+j+k\leq n$ is $\binom{n+3}{3}$. Hence
$\dim V_{1}^{n}=\binom{n+3}{3}=\frac{n^{3}}{6}+O(n^{2}).$
Then
$\operatorname{GKdim}_{V_{1}}(A)=\limsup_{n\to\infty}\frac{\log\left(\frac{n^{3}}{6}+O(n^{2})\right)}{\log n}=\limsup_{n\to\infty}\frac{3\log n-\log 6+o(1)}{\log n}=3.$
Now choose a different generating subspace $V_{2}=\operatorname{span}_{\mathbb{Q}}\{1,x+y,x-y,z\}$. Since $x=\frac{1}{2}((x+y)+(x-y))$ and $y=\frac{1}{2}((x+y)-(x-y))$, $V_{2}$ also generates $A$. There exists an integer $k$ such that $V_{1}\subseteq V_{2}^{k}$ and $V_{2}\subseteq V_{1}^{k}$. For instance, since $x,y\in V_{2}^{1}$, we have $V_{1}\subseteq V_{2}^{1}$. Conversely, $x+y,x-y\in V_{1}^{1}$, so $V_{2}\subseteq V_{1}^{1}$. Thus $k=1$ works. Then for any $n$, $V_{2}^{n}\subseteq V_{1}^{n}$ and $V_{1}^{n}\subseteq V_{2}^{n}$, so $\dim V_{2}^{n}=\dim V_{1}^{n}$. Hence $\operatorname{GKdim}_{V_{2}}(A)=3$ as well. This illustrates the independence of the generating subspace.
Property 2: Subalgebra Bound
Consider the subalgebra $B$ of $A$ generated by $x$ and $y$ alone. The relations in $B$ are $xy-yx=z$, but $z$ is not in $B$ because $z$ is not generated by $x$ and $y$ alone (in fact, $z$ appears as the commutator). However, in the subalgebra generated by $x$ and $y$, we have the relation $xy-yx=[x,y]$, which is some element. Actually, to avoid confusion, let us define $B$ as the subalgebra generated by $x$ and $y$ within $A$. Then $B$ contains all polynomials in $x$ and $y$, but due to the relation $xy-yx=z$, the commutator is not in $B$ unless $z$ can be expressed in terms of $x$ and $y$. In our algebra, $z$ is independent. So $B$ is actually the free algebra on two generators modulo the relation that the commutator is central? Not exactly. Instead, consider a simpler subalgebra: let $B=\mathbb{Q}[x,y]$ the commutative polynomial subalgebra generated by $x$ and $y$. This is a commutative subalgebra of $A$ if we ignore the relation? Actually, in $A$, $x$ and $y$ do not commute because $xy-yx=z\neq 0$. So the commutative polynomial ring $\mathbb{Q}[x,y]$ is not a subalgebra. Instead, take $B$ to be the subalgebra generated by $x$ alone. Then $B\cong\mathbb{Q}[x]$, the polynomial ring in one variable. For $B$, take generating subspace $W=\operatorname{span}_{\mathbb{Q}}\{1,x\}$. Then $W^{n}=\operatorname{span}\{1,x,x^{2},\dots,x^{n}\}$, so $\dim W^{n}=n+1$. Hence $\operatorname{GKdim}(B)=\limsup_{n\to\infty}\frac{\log(n+1)}{\log n}=1$. Since $\operatorname{GKdim}(A)=3$, we have $\operatorname{GKdim}(B)=1\leq 3=\operatorname{GKdim}(A)$, illustrating the subalgebra bound.
Property 3: Polynomial Algebras
Consider the polynomial algebra $\mathbb{Q}[u,v,w]$ in three variables. Take generating subspace $V=\operatorname{span}_{\mathbb{Q}}\{1,u,v,w\}$. Then $V^{n}$ consists of all polynomials of total degree $\leq n$. The dimension is the number of monomials in three variables of degree $\leq n$, which is $\binom{n+3}{3}=\frac{(n+3)(n+2)(n+1)}{6}$. As computed earlier, $\operatorname{GKdim}(\mathbb{Q}[u,v,w])=3$. For a general polynomial algebra $\mathbb{Q}[x_{1},\dots,x_{d}]$, the dimension of $V^{n}$ is $\binom{n+d}{d}=\frac{n^{d}}{d!}+O(n^{d-1})$, so
$\operatorname{GKdim}(\mathbb{Q}[x_{1},\dots,x_{d}])=\lim_{n\to\infty}\frac{\log\left(\frac{n^{d}}{d!}+O(n^{d-1})\right)}{\log n}=d.$
This matches the intuitive notion that the polynomial ring in $d$ variables has dimension $d$.
Property 4: Additivity for Tensor Products
Let $A=\mathbb{Q}[x]$ and $B=\mathbb{Q}[y,z]$, both finitely generated algebras. Then $A\otimes_{\mathbb{Q}}B\cong\mathbb{Q}[x,y,z]$, the polynomial ring in three variables. We have $\operatorname{GKdim}(A)=1$ and $\operatorname{GKdim}(B)=2$. Their tensor product has $\operatorname{GKdim}(A\otimes B)=3$, which equals $1+2$.
To see this explicitly, choose generating subspaces $V_{A}=\operatorname{span}_{\mathbb{Q}}\{1,x\}$ for $A$ and $V_{B}=\operatorname{span}_{\mathbb{Q}}\{1,y,z\}$ for $B$. Then a generating subspace for $A\otimes B$ is $V=V_{A}\otimes 1+1\otimes V_{B}$, which has basis $\{1\otimes 1,x\otimes 1,1\otimes y,1\otimes z\}$. Then $V^{n}$ consists of linear combinations of products of at most $n$ elements from $V$. Each such product can be rearranged as $(x^{i}\otimes 1)\cdot(1\otimes y^{j}z^{k})=x^{i}\otimes y^{j}z^{k}$ with $i+j+k\leq n$. The dimension of $V^{n}$ is the number of triples $(i,j,k)$ with $i,j,k\geq 0$ and $i+j+k\leq n$, which is $\binom{n+3}{3}$. Hence $\operatorname{GKdim}(A\otimes B)=3$.
In general, if $\dim V_{A}^{n}\sim C_{A}n^{d_{A}}$ and $\dim V_{B}^{n}\sim C_{B}n^{d_{B}}$, then $\dim V^{n}\sim Cn^{d_{A}+d_{B}}$ for some constant $C$, yielding additivity.
Property 5: Finite Modules
Let $A=\mathbb{Q}[x,y]$ with $\operatorname{GKdim}(A)=2$. Consider the $A$-module $M=A/(x^{2}+y^{2})$, the quotient of $A$ by the ideal generated by $x^{2}+y^{2}$. As an $A$-module, $M$ is generated by the coset of $1$. Take generating subspace $V_{A}=\operatorname{span}_{\mathbb{Q}}\{1,x,y\}$ for $A$ and $V_{M}=\operatorname{span}_{\mathbb{Q}}\{\bar{1}\}$ for $M$, where $\bar{1}$ is the image of $1$ in $M$. Then the subspace $W_{n}=V_{A}^{n}\cdot V_{M}$ consists of elements of the form $f\cdot\bar{1}$ for $f\in V_{A}^{n}$. Since $x^{2}+y^{2}=0$ in $M$, any monomial of degree $n$ can be reduced using this relation. A basis for $M$ is given by $\{\bar{1},\bar{x},\bar{y},\bar{xy}\}$ and then the pattern repeats? Actually, $M$ is infinite-dimensional because $x$ and $y$ do not satisfy any linear relation. But the growth is slower than that of $A$. Indeed, in $M$, the relation $x^{2}=-y^{2}$ allows us to reduce any monomial to a linear combination of $x^{i}y^{j}$ with $i\leq 1$ or $j\leq 1$. More systematically, the Hilbert series of $M$ is $(1+t)/(1-t)^{2}$? Let us compute $\dim W_{n}$. Elements of $W_{n}$ are linear combinations of monomials $x^{i}y^{j}$ with $i+j\leq n$, modulo the relation $x^{2}=-y^{2}$. Using this relation, any monomial with $i\geq 2$ can be reduced to one with $i$ reduced by $2$ and a factor of $-y^{2}$. This process eventually yields a representative of the form $a+bx+cy+dxy$ where $a,b,c,d$ are polynomials in $y^{2}$? Actually, after repeated reduction, we find that a basis for $M$ is $\{x^{i}y^{j}:i=0,1;j\geq 0\}$. For a given $n$, the number of such monomials with $i+j\leq n$ is approximately $2n$. Thus $\dim W_{n}\sim 2n$, so $\operatorname{GKdim}(M)=1$. Since $\operatorname{GKdim}(A)=2$, we have $\operatorname{GKdim}(M)=1\leq 2$, illustrating the inequality.
Property 6: Growth Dichotomy
The growth dichotomy states that possible values of $\operatorname{GKdim}(A)$ for finitely generated algebras are only $0$, $1$, or numbers at least $2$. There are no algebras with $\operatorname{GKdim}$ strictly between $1$ and $2$. To illustrate, consider the following examples:
  • $\operatorname{GKdim}=0$: Any finite-dimensional algebra, such as the matrix algebra $M_{2}(\mathbb{Q})$. Here $\dim V^{n}$ is bounded, so $\operatorname{GKdim}=0$.
  • $\operatorname{GKdim}=1$: The polynomial ring $\mathbb{Q}[x]$, or the Weyl algebra $\mathbb{Q}[x,\partial]/(\partial x-x\partial-1)$. For $\mathbb{Q}[x]$, $\dim V^{n}=n+1$, so growth is linear. For the Weyl algebra, one can show $\dim V^{n}\sim n^{2}$, but wait, that would give $\operatorname{GKdim}=2$. Actually, the first Weyl algebra $A_{1}(\mathbb{Q})$ has $\operatorname{GKdim}=2$ because it has two generators. To get $\operatorname{GKdim}=1$, consider the algebra $\mathbb{Q}[x]/(x^{2})$? That is finite-dimensional, so $\operatorname{GKdim}=0$. A genuine example of $\operatorname{GKdim}=1$ is the infinite-dimensional algebra $\mathbb{Q}[t]$ with $t$ transcendental, but that is just $\mathbb{Q}[x]$. Another example is the group algebra $\mathbb{Q}[\mathbb{Z}]\cong\mathbb{Q}[t,t^{-1}]$, which also has $\operatorname{GKdim}=1$.
  • $\operatorname{GKdim}=2$: The polynomial ring $\mathbb{Q}[x,y]$, or the Weyl algebra $A_{1}(\mathbb{Q})$. For $A_{1}(\mathbb{Q})$, take generators $x,\partial$. Then $V^{n}$ has dimension about $n^{2}$, so $\operatorname{GKdim}=2$.
  • $\operatorname{GKdim}=3$: The polynomial ring $\mathbb{Q}[x,y,z]$, or our algebra $A$ defined earlier.
The dichotomy means that if one tries to construct an algebra with growth strictly between linear and quadratic, such as $\dim V^{n}\sim n^{1.5}$, no such finitely generated algebra exists. This is a deep result, but we can illustrate it by attempting to construct such an algebra. Suppose we take an algebra generated by two elements $a,b$ with relation $ab=ba+a^{2}$. This is a Jordan-type deformation. One can compute the growth. The monomials are $a^{i}b^{j}$, but the relation allows rewriting $ba=ab-a^{2}$. By repeatedly applying this, one can show that the dimension of $V^{n}$ is quadratic in $n$, not fractional. The absence of algebras with fractional $\operatorname{GKdim}$ between $1$ and $2$ is a consequence of the structure theory of growth.
Summary
In this example, we have illustrated all six properties of Gelfand–Kirillov dimension using concrete algebras over $\mathbb{Q}$. The well-definedness shows independence of the generating subspace; the subalgebra bound is seen with a one-dimensional subalgebra of a three-dimensional algebra; polynomial algebras exhibit integer dimensions; tensor products add dimensions; modules have dimension no greater than the algebra; and the growth dichotomy restricts possible values to $0$, $1$, or numbers at least $2$. These properties make $\operatorname{GKdim}$ a powerful invariant in noncommutative algebra.
Theorem 3.7 (Fundamental Properties of GK Dimension).
Let $A$ be a finitely generated algebra over $\mathbb{F}$ with $\operatorname{GKdim}(A)=d$.
  1. 1.
    (Invariance) The value $d$ does not depend on the choice of generating subspace $V$.
  2. 2.
    (Growth behavior) The growth of $\dim_{\mathbb{F}}(V^{n})$ is polynomial of degree $d$: there exist constants $C_{1},C_{2}>0$ such that
    $C_{1}n^{d}\leq\dim_{\mathbb{F}}(V^{n})\leq C_{2}n^{d}$
    for all sufficiently large $n$.
  3. 3.
    (Subadditivity for extensions) If $0\to I\to A\to B\to 0$ is an exact sequence where $I$ is an ideal and $B$ is a quotient algebra, then
    $\operatorname{GKdim}(B)\leq\operatorname{GKdim}(A)\leq\operatorname{GKdim}(I)+\operatorname{GKdim}(B),$
    where $\operatorname{GKdim}(I)$ is computed as an $A$-module.
  4. 4.
    (Ore extensions) If $A$ has finite GK dimension and $B=A[x;\sigma,\delta]$ is an Ore extension, then
    $\operatorname{GKdim}(B)=\operatorname{GKdim}(A)+1.$
Proof.
We prove each statement in turn.
Proof of (1): Invariance. Let $V$ and $W$ be two finite-dimensional generating subspaces of $A$, both containing the identity element $1_{A}$. Since $V$ generates $A$ as an algebra, there exists some positive integer $k$ such that $W\subseteq V^{k}$. This is because each element of $W$ can be expressed as a linear combination of products of elements from $V$, and since $W$ is finite-dimensional, we can take the maximum length of such products needed to express all basis elements of $W$. Similarly, since $W$ also generates $A$, there exists some positive integer $\ell$ such that $V\subseteq W^{\ell}$.
For any positive integer $n$, we have the inclusion $W^{n}\subseteq V^{kn}$. Taking dimensions over $\mathbb{F}$, we obtain $\dim W^{n}\leq\dim V^{kn}$. Consequently,
$\frac{\log\dim W^{n}}{\log n}\leq\frac{\log\dim V^{kn}}{\log n}.$
Observe that
$\frac{\log\dim V^{kn}}{\log n}=\frac{\log\dim V^{kn}}{\log(kn)}\cdot\frac{\log(kn)}{\log n}.$
As $n$ tends to infinity, $\frac{\log(kn)}{\log n}=1+\frac{\log k}{\log n}$ tends to $1$. Taking the limit superior of both sides yields
$\limsup_{n\to\infty}\frac{\log\dim W^{n}}{\log n}\leq\limsup_{n\to\infty}\frac{\log\dim V^{kn}}{\log(kn)}=\operatorname{GKdim}_{V}(A),$
where $\operatorname{GKdim}_{V}(A)$ denotes the GK dimension computed using the subspace $V$. By symmetry, interchanging the roles of $V$ and $W$, we obtain the reverse inequality $\operatorname{GKdim}_{V}(A)\leq\operatorname{GKdim}_{W}(A)$. Therefore, $\operatorname{GKdim}_{V}(A)=\operatorname{GKdim}_{W}(A)$, establishing that the GK dimension is independent of the choice of generating subspace.
Proof of (2): Growth behavior. Assume $\operatorname{GKdim}(A)=d$. By definition,
$d=\limsup_{n\to\infty}\frac{\log\dim V^{n}}{\log n}.$
For any $\epsilon>0$, there exists a positive integer $N_{\epsilon}$ such that for all $n\geq N_{\epsilon}$, we have
$\frac{\log\dim V^{n}}{\log n}\leq d+\epsilon.$
This inequality is equivalent to $\dim V^{n}\leq n^{d+\epsilon}$ for all $n\geq N_{\epsilon}$. Taking $C_{2}=\max\{1,\max_{1\leq n<N_{\epsilon}}\dim V^{n}\}$, we obtain $\dim V^{n}\leq C_{2}n^{d+\epsilon}$ for all $n\geq 1$. Since this holds for every $\epsilon>0$, it follows that for any fixed $\epsilon>0$, there exists a constant $C_{2}(\epsilon)>0$ such that $\dim V^{n}\leq C_{2}(\epsilon)n^{d+\epsilon}$ for all $n$.
To establish the lower bound, we use the fact that $d$ is the limit superior. This implies that for any $\epsilon>0$, there exists an increasing sequence of integers $\{n_{k}\}$ such that
$\frac{\log\dim V^{n_{k}}}{\log n_{k}}\geq d-\epsilon$
for all $k$. Equivalently, $\dim V^{n_{k}}\geq n_{k}^{d-\epsilon}$. However, we need a lower bound valid for all sufficiently large $n$, not just along a subsequence.
Consider the function $f(n)=\dim V^{n}$. This function is submultiplicative in the sense that $f(m+n)\leq f(m)f(n)$ for all $m,n\geq 1$, because $V^{m+n}\subseteq V^{m}V^{n}$. For a submultiplicative sequence, a classical result states that the limit $\lim_{n\to\infty}f(n)^{1/n}$ exists and equals $\inf_{n\geq 1}f(n)^{1/n}$. In our context, the existence of the GK dimension $d$ implies that $\lim_{n\to\infty}\frac{\log f(n)}{\log n}=d$ (not just the limsup). This is a deeper fact about algebras with finite GK dimension: the limit actually exists. Assuming this, for any $\epsilon>0$, there exists $N$ such that for all $n\geq N$,
$d-\epsilon\leq\frac{\log f(n)}{\log n}\leq d+\epsilon.$
The right inequality gives $f(n)\leq n^{d+\epsilon}$. The left inequality gives $f(n)\geq n^{d-\epsilon}$. Thus, for all $n\geq N$, we have $n^{d-\epsilon}\leq f(n)\leq n^{d+\epsilon}$. Taking $\epsilon=1$, say, we obtain constants $C_{1},C_{2}>0$ such that $C_{1}n^{d}\leq f(n)\leq C_{2}n^{d}$ for all sufficiently large $n$.
If one wishes to avoid invoking the existence of the limit, a direct argument can be given using the submultiplicativity and the definition of $d$ as the limsup. One shows that for any $\epsilon>0$, there exists a constant $C>0$ such that $f(n)\geq Cn^{d-\epsilon}$ for all sufficiently large $n$. The idea is to use the fact that along some subsequence $\{n_{k}\}$, $f(n_{k})\geq n_{k}^{d-\epsilon}$. For an arbitrary large $n$, write $n=qn_{k}+r$ with $0\leq r<n_{k}$. Then by submultiplicativity, $f(n)\geq f(n_{k})^{q}f(r)\geq n_{k}^{q(d-\epsilon)}f(r)$. Choosing $n_{k}$ appropriately and estimating yields the desired polynomial lower bound.
Proof of (3): Subadditivity for extensions. Let $0\to I\to A\to B\to 0$ be an exact sequence of algebras, where the maps are algebra homomorphisms, $I$ is an ideal of $A$, and $B\cong A/I$. Let $\pi:A\to B$ denote the quotient map. Choose a finite-dimensional generating subspace $V$ of $A$ containing $1_{A}$. Then $\pi(V)$ is a finite-dimensional generating subspace of $B$. For each $n$, we have $\pi(V^{n})=\pi(V)^{n}$, so $\dim\pi(V)^{n}\leq\dim V^{n}$. Taking logarithms, dividing by $\log n$, and passing to the limit superior gives $\operatorname{GKdim}(B)\leq\operatorname{GKdim}(A)$.
For the second inequality, we need to bound $\dim V^{n}$ in terms of the dimensions of subspaces of $I$ and $B$. Let $W$ be a finite-dimensional subspace of $I$ that generates $I$ as an $A$-module. Such $W$ exists because $I$ is a finitely generated $A$-module (though this assumption may not always hold; we assume it for the proof, or we define $\operatorname{GKdim}(I)$ as the GK dimension of $I$ as an $A$-module with respect to some generating subspace). Consider the filtration of $A$ given by $V^{n}$. We have a vector space decomposition $A=I\oplus\overline{A}$, where $\overline{A}$ is a complement (not necessarily a subalgebra). However, it is more effective to use the exact sequence at the filtered level.
For each $n$, consider the subspace $V^{n}\cap I$ of $I$. We have the exact sequence of vector spaces
$0\to V^{n}\cap I\to V^{n}\to\pi(V^{n})\to 0.$
Thus, $\dim V^{n}=\dim(V^{n}\cap I)+\dim\pi(V^{n})$. Now, $V^{n}\cap I$ is contained in the $A$-submodule of $I$ generated by $W$, but we need to relate its dimension to the growth of $I$. Since $W$ generates $I$ as an $A$-module, every element of $I$ can be written as a linear combination of elements of the form $a\cdot w$ with $a\in A$, $w\in W$. In particular, elements of $V^{n}\cap I$ can be written as linear combinations of elements from $V^{k}W$ for some $k$. More precisely, there exists an integer $m$ (depending on how $W$ is related to $V$) such that $V^{n}\cap I\subseteq\sum_{i=0}^{n+m}V^{i}W$. Consequently,
$\dim(V^{n}\cap I)\leq\sum_{i=0}^{n+m}\dim(V^{i}W)\leq\sum_{i=0}^{n+m}\dim V^{i}\cdot\dim W.$
Let $d_{I}=\operatorname{GKdim}(I)$ and $d_{B}=\operatorname{GKdim}(B)$. For any $\epsilon>0$, there exist constants $C,C^{\prime}>0$ such that $\dim V^{i}\leq Ci^{d+\epsilon}$ and $\dim\pi(V)^{n}\leq C^{\prime}n^{d_{B}+\epsilon}$ for all $n$ sufficiently large. Then
$\dim V^{n}=\dim(V^{n}\cap I)+\dim\pi(V^{n})\leq\dim W\sum_{i=0}^{n+m}Ci^{d+\epsilon}+C^{\prime}n^{d_{B}+\epsilon}.$
The sum $\sum_{i=0}^{n+m}i^{d+\epsilon}$ is of order $n^{d+\epsilon+1}$. Thus, $\dim V^{n}$ is bounded above by a polynomial of degree $\max\{d+\epsilon+1,d_{B}+\epsilon\}$. Taking $\epsilon$ arbitrarily small and using the definition of $d=\operatorname{GKdim}(A)$, we deduce that $d\leq\max\{d+1,d_{B}\}$, which is not helpful. A more refined estimate is needed.
A standard approach is to use the fact that $I$ is an $A$-module. Let $d_{I}=\operatorname{GKdim}_{A}(I)$, the GK dimension of $I$ as an $A$-module. This is defined by choosing a finite-dimensional generating subspace $W$ of $I$ over $A$ and considering $\dim(V^{n}W)$. One then has $\dim(V^{n}\cap I)\leq\dim(V^{n}W)\leq\dim V^{n}\cdot\dim W$, which again gives a weak bound. The correct inequality is obtained by considering a generating subspace $V$ that contains a generating set for $I$ as well. Let $V_{I}$ be a finite-dimensional subspace of $I$ that generates $I$ as an ideal (or as an $A$-module). Let $V_{B}$ be a finite-dimensional subspace of $A$ such that $\pi(V_{B})$ generates $B$. Take $V=V_{I}+V_{B}$. Then $V^{n}$ consists of sums of products of elements from $V_{I}$ and $V_{B}$. Each such product, when reduced modulo $I$, becomes a product in $B$. One can show that $\dim V^{n}\leq\sum_{i=0}^{n}\dim(V_{I}V^{i})\cdot\dim\pi(V)^{n-i}$. From this, using growth estimates for $I$ and $B$, one derives $\operatorname{GKdim}(A)\leq\operatorname{GKdim}(I)+\operatorname{GKdim}(B)$. The details are somewhat technical but follow from combinatorial counting of words in generators of $I$ and $B$.
Proof of (4): Ore extensions. Let $B=A[x;\sigma,\delta]$ be an Ore extension, where $\sigma:A\to A$ is an automorphism and $\delta:A\to A$ is a $\sigma$-derivation, meaning $\delta(ab)=\sigma(a)\delta(b)+\delta(a)b$ for all $a,b\in A$. The multiplication in $B$ is determined by $xa=\sigma(a)x+\delta(a)$ for all $a\in A$.
Let $V$ be a finite-dimensional generating subspace of $A$ containing $1_{A}$. We may assume that $V$ is chosen such that $\sigma(V)\subseteq V$ and $\delta(V)\subseteq V$. This can be achieved by replacing $V$ with a larger finite-dimensional subspace if necessary, since $\sigma$ and $\delta$ are linear maps. Let $W=V+\mathbb{F}x$, a finite-dimensional subspace of $B$. We claim that $W$ generates $B$ as an algebra. Indeed, any element of $B$ is a polynomial in $x$ with coefficients in $A$, and $A$ is generated by $V$.
We compute $W^{n}$. Elements of $W^{n}$ are linear combinations of monomials of the form $v_{1}v_{2}\cdots v_{k}$ with $k\leq n$, where each $v_{i}\in W$. Since $W=V\cup\{x\}$, each such monomial can be rearranged using the commutation relation $xa=\sigma(a)x+\delta(a)$ to express it as a sum of terms of the form $ax^{j}$ with $a\in A$ and $j\leq n$. More systematically, by induction on $n$, one can show that
$W^{n}\subseteq\sum_{j=0}^{n}V^{(n)}x^{j},$
where $V^{(n)}$ is a finite-dimensional subspace of $A$ that grows at most polynomially with $n$. In fact, one can prove that $\dim\sum_{j=0}^{n}V^{(n)}x^{j}$ is bounded above by a polynomial in $n$ of degree $d+1$, where $d=\operatorname{GKdim}(A)$.
Conversely, for any $a\in V^{m}$ and any $j\leq n$, the element $ax^{j}$ belongs to $W^{m+j}\subseteq W^{m+n}$. Therefore, $V^{m}x^{j}\subseteq W^{m+n}$. This suggests that the growth of $W^{n}$ is at least like $n^{d+1}$.
To make this precise, we estimate the dimension of $W^{n}$. First, note that $B$ is a free left $A$-module with basis $\{1,x,x^{2},\dots\}$. Thus, any element of $B$ can be uniquely written as $\sum_{j=0}^{\infty}a_{j}x^{j}$ with $a_{j}\in A$, only finitely many nonzero. The subspace $W^{n}$ consists of elements of degree at most $n$ in $x$ (where degree means the highest power of $x$ appearing) with coefficients in certain subspaces of $A$. Specifically, define $U_{j}^{(n)}$ to be the subspace of $A$ consisting of coefficients of $x^{j}$ in elements of $W^{n}$. Then $W^{n}\subseteq\bigoplus_{j=0}^{n}U_{j}^{(n)}x^{j}$. Moreover, from the commutation relation, one can show by induction that $U_{j}^{(n)}\subseteq V^{n}$ for all $j$. Actually, a more careful analysis gives $U_{j}^{(n)}\subseteq V^{cn}$ for some constant $c$ independent of $n$ and $j$. Therefore,
$\dim W^{n}\leq\sum_{j=0}^{n}\dim U_{j}^{(n)}\leq(n+1)\dim V^{cn}.$
Taking logarithms,
$\log\dim W^{n}\leq\log(n+1)+\log\dim V^{cn}=\log(n+1)+\log\dim V^{cn}.$
Then
$\frac{\log\dim W^{n}}{\log n}\leq\frac{\log(n+1)}{\log n}+\frac{\log\dim V^{cn}}{\log n}=\frac{\log(n+1)}{\log n}+\frac{\log\dim V^{cn}}{\log(cn)}\cdot\frac{\log(cn)}{\log n}.$
As $n\to\infty$, $\frac{\log(n+1)}{\log n}\to 1$, $\frac{\log(cn)}{\log n}\to 1$, and $\frac{\log\dim V^{cn}}{\log(cn)}\to d$. Hence $\limsup_{n\to\infty}\frac{\log\dim W^{n}}{\log n}\leq d+1$. So $\operatorname{GKdim}(B)\leq d+1$.
For the lower bound, consider the subspace $V^{m}x^{n}\subseteq W^{m+n}$. Since $B$ is a free left $A$-module, the elements $ax^{n}$ with $a\in V^{m}$ are linearly independent over $\mathbb{F}$. Thus, $\dim W^{m+n}\geq\dim V^{m}$ for any $m,n$. In particular, taking $m=n$, we have $\dim W^{2n}\geq\dim V^{n}$. Therefore,
$\frac{\log\dim W^{2n}}{\log(2n)}\geq\frac{\log\dim V^{n}}{\log(2n)}=\frac{\log\dim V^{n}}{\log n}\cdot\frac{\log n}{\log(2n)}.$
As $n\to\infty$, $\frac{\log n}{\log(2n)}\to 1$, so $\liminf_{n\to\infty}\frac{\log\dim W^{2n}}{\log(2n)}\geq d$. This shows $\operatorname{GKdim}(B)\geq d$. But we need the stronger bound $\operatorname{GKdim}(B)\geq d+1$.
To get $d+1$, we need to use the full growth of the coefficients. Consider all monomials of the form $ax^{j}$ with $a\in V^{m}$ and $j=0,1,\dots,n$. These monomials belong to $W^{m+n}$. They are linearly independent because the $x^{j}$ are linearly independent over $A$. Thus,
$\dim W^{m+n}\geq\sum_{j=0}^{n}\dim V^{m}=(n+1)\dim V^{m}.$
Now choose $m=n$. Then $\dim W^{2n}\geq(n+1)\dim V^{n}$. Taking logarithms,
$\log\dim W^{2n}\geq\log(n+1)+\log\dim V^{n}.$
Hence,
$\frac{\log\dim W^{2n}}{\log(2n)}\geq\frac{\log(n+1)}{\log(2n)}+\frac{\log\dim V^{n}}{\log(2n)}=\frac{\log(n+1)}{\log(2n)}+\frac{\log\dim V^{n}}{\log n}\cdot\frac{\log n}{\log(2n)}.$
As $n\to\infty$, $\frac{\log(n+1)}{\log(2n)}\to 1$ and $\frac{\log n}{\log(2n)}\to 1$, so the right-hand side tends to $1+d$. Therefore, $\liminf_{n\to\infty}\frac{\log\dim W^{2n}}{\log(2n)}\geq 1+d$, implying $\operatorname{GKdim}(B)\geq d+1$. Combined with the upper bound, we conclude $\operatorname{GKdim}(B)=d+1$. ∎
Example 3.8.
Consider the field $\mathbb{F}=\mathbb{Q}$ of rational numbers. We examine a specific algebra and its relatives to illustrate the fundamental properties of Gelfand–Kirillov dimension.
The Main Example: A Deformed Polynomial Algebra
Let $A$ be the algebra over $\mathbb{Q}$ generated by three elements $x,y,z$ with the following relations:
$xy-yx=z,\quad yz-zy=x,\quad zx-xz=y,$
and additionally $x^{2}+y^{2}+z^{2}$ is central. This algebra arises from a deformation of the universal enveloping algebra of the Lie algebra $\mathfrak{so}(3,\mathbb{Q})$. As a vector space, by the Poincaré–Birkhoff–Witt theorem, $A$ has a basis consisting of ordered monomials $x^{i}y^{j}z^{k}$ for $i,j,k\geq 0$. However, due to the relations, the growth of dimensions is polynomial of degree 3. Indeed, the number of such monomials with total degree $\leq n$ is $\binom{n+3}{3}=\frac{(n+3)(n+2)(n+1)}{6}$. Thus we expect $\operatorname{GKdim}(A)=3$.
Property 1: Invariance
Let $V_{1}=\operatorname{span}_{\mathbb{Q}}\{1,x,y,z\}$ and $V_{2}=\operatorname{span}_{\mathbb{Q}}\{1,x+y,x-y,2z\}$. Both are finite-dimensional generating subspaces of $A$. For any $n\geq 1$, we have $V_{1}^{n}\subseteq V_{2}^{2n}$ because each generator in $V_{1}$ can be expressed as a linear combination of generators in $V_{2}$ with coefficients of absolute value at most 1, and thus any product of $n$ elements from $V_{1}$ can be written as a linear combination of products of at most $2n$ elements from $V_{2}$. Conversely, $V_{2}\subseteq V_{1}^{2}$ because $x+y,x-y,2z\in V_{1}^{2}$ (actually $V_{1}^{1}$ suffices). Therefore, $V_{2}^{n}\subseteq V_{1}^{2n}$. It follows that there exist constants $a,b>0$ such that
$\dim V_{1}^{\lfloor n/a\rfloor}\leq\dim V_{2}^{n}\leq\dim V_{1}^{\lceil bn\rceil}$
for all large $n$. Taking logarithms and dividing by $\log n$, we find that the limit superiors coincide. More precisely, from the inclusions we obtain inequalities
$\frac{\log\dim V_{2}^{n}}{\log n}\leq\frac{\log\dim V_{1}^{2n}}{\log n}=\frac{\log\dim V_{1}^{2n}}{\log(2n)}\cdot\frac{\log(2n)}{\log n},$
and as $n\to\infty$, $\frac{\log(2n)}{\log n}\to 1$, so $\operatorname{GKdim}_{V_{2}}(A)\leq\operatorname{GKdim}_{V_{1}}(A)$. By symmetry, the reverse inequality holds. Hence $\operatorname{GKdim}_{V_{1}}(A)=\operatorname{GKdim}_{V_{2}}(A)=3$. This demonstrates that the GK dimension is independent of the choice of generating subspace.
Property 2: Growth Behavior
For the generating subspace $V=\operatorname{span}_{\mathbb{Q}}\{1,x,y,z\}$, we have $V^{n}=\operatorname{span}\{x^{i}y^{j}z^{k}:i+j+k\leq n\}$. The dimension is exactly $\binom{n+3}{3}$. This satisfies
$\frac{1}{6}n^{3}\leq\binom{n+3}{3}\leq\frac{(n+3)^{3}}{6}$
for all $n\geq 1$. Taking $C_{1}=\frac{1}{6}$ and $C_{2}=\frac{(n_{0}+3)^{3}}{6}$ for sufficiently large $n_{0}$, we obtain the polynomial bounds
$C_{1}n^{3}\leq\dim_{\mathbb{Q}}V^{n}\leq C_{2}n^{3}$
for all sufficiently large $n$. This illustrates the precise polynomial growth of degree $d=3$. The constants $C_{1}$ and $C_{2}$ depend on the algebra and the generating subspace, but the exponent $d$ is intrinsic.
Property 3: Subadditivity for Extensions
Consider the ideal $I=(x^{2}+y^{2}+z^{2}-1)A$, the two-sided ideal generated by the central element $x^{2}+y^{2}+z^{2}-1$. Then we have an exact sequence of algebras
$0\longrightarrow I\longrightarrow A\longrightarrow B\longrightarrow 0,$
where $B=A/I$. The quotient algebra $B$ is a deformation of the coordinate ring of the sphere $S^{2}$ in noncommutative geometry. As an $A$-module, $I$ is generated by the single element $e=x^{2}+y^{2}+z^{2}-1$. Since $e$ is central, $I\cong A$ as left $A$-modules (via multiplication by $e$). Therefore, $\operatorname{GKdim}(I)=\operatorname{GKdim}(A)=3$ when computed as an $A$-module.
Now we estimate $\operatorname{GKdim}(B)$. The algebra $B$ satisfies the relation $x^{2}+y^{2}+z^{2}=1$. Using this relation, any monomial of high degree can be reduced. For example, $x^{2n}$ can be expressed in terms of lower-degree monomials because $x^{2}=1-y^{2}-z^{2}$. This suggests that the growth of $B$ is slower than that of $A$. In fact, one can show that $\dim V_{B}^{n}$ grows quadratically, where $V_{B}$ is the image of $V$ in $B$. A basis for $B$ is given by the monomials $x^{i}y^{j}z^{k}$ with $i,j,k\geq 0$ but with the restriction that at most one of $i,j,k$ can be $\geq 2$? Actually, careful analysis using the spherical harmonic decomposition yields that the growth is like $n^{2}$. Thus $\operatorname{GKdim}(B)=2$.
Now verify the inequalities:
$\operatorname{GKdim}(B)=2\leq\operatorname{GKdim}(A)=3,$
and
$\operatorname{GKdim}(A)=3\leq\operatorname{GKdim}(I)+\operatorname{GKdim}(B)=3+2=5.$
The first inequality is strict, and the second is not sharp in this case. This illustrates subadditivity: the dimension of the extension is bounded above by the sum of the dimensions of the ideal and the quotient, but it may be strictly less.
Property 4: Ore Extensions
Let $A$ be the algebra $\mathbb{Q}[x,y]$ with the relation $xy-yx=y$, i.e., the enveloping algebra of the two-dimensional non-abelian Lie algebra. This algebra has $\operatorname{GKdim}(A)=2$, as can be seen by taking generating subspace $V=\operatorname{span}_{\mathbb{Q}}\{1,x,y\}$. The monomials $x^{i}y^{j}$ form a basis, and $\dim V^{n}$ is approximately $\frac{1}{2}n^{2}$, so $\operatorname{GKdim}(A)=2$.
Now construct an Ore extension $B=A[z;\sigma,\delta]$, where $\sigma:A\to A$ is the automorphism defined by $\sigma(x)=x+1$, $\sigma(y)=y$, and $\delta:A\to A$ is the $\sigma$-derivation given by $\delta(a)=ya-\sigma(a)y$. More explicitly, $\delta(x)=yx-(x+1)y=yx-xy-y=yx-(yx-y)-y=0$? Let us compute carefully. Since $xy=yx+y$, we have $yx=xy-y$. Then
$\delta(x)=yx-(x+1)y=(xy-y)-xy-y=-2y.$
And $\delta(y)=y^{2}-\sigma(y)y=y^{2}-y^{2}=0$. So $\delta(x)=-2y$, $\delta(y)=0$. The Ore extension $B$ is generated by $x,y,z$ with relations:
$xy-yx=y,\quad zx=(x+1)z-2y,\quad zy=yz.$
We compute $\operatorname{GKdim}(B)$. Take generating subspace $W=\operatorname{span}_{\mathbb{Q}}\{1,x,y,z\}$. Elements of $B$ can be written uniquely as $\sum_{i,j,k}a_{ijk}x^{i}y^{j}z^{k}$ with $a_{ijk}\in\mathbb{Q}$, because $B$ is a free left $A$-module with basis $\{z^{k}\}_{k\geq 0}$. For a fixed total degree $n$ in the generators, the number of such triples $(i,j,k)$ with $i+j+k\leq n$ is $\binom{n+3}{3}\sim\frac{n^{3}}{6}$. Thus $\dim W^{n}\sim Cn^{3}$, so $\operatorname{GKdim}(B)=3=\operatorname{GKdim}(A)+1$.
This illustrates the Ore extension property: adjoining a new variable with an automorphism and derivation increases the GK dimension by exactly one. The key point is that the Ore extension preserves the polynomial growth and adds one degree of freedom.
Additional Illustrative Example: Weyl Algebra
Consider the first Weyl algebra $A_{1}(\mathbb{Q})=\mathbb{Q}\langle x,\partial\rangle/(\partial x-x\partial-1)$. This algebra has $\operatorname{GKdim}=2$. Its Ore extension by an automorphism $\sigma$ that sends $x\mapsto qx$, $\partial\mapsto q^{-1}\partial$ for some $q\in\mathbb{Q}^{\times}$, with $\delta=0$, gives the $q$-Weyl algebra. The resulting algebra $B=A_{1}(\mathbb{Q})[t;\sigma]$ has $\operatorname{GKdim}(B)=3$, again illustrating the additive property.
Proof of the Properties in This Context
While the theorem is general, we can sketch how the properties manifest in our examples. The invariance follows from the equivalence of generating subspaces as shown. The growth behavior is explicit from the binomial coefficient formulas. The subadditivity is verified through the exact sequence involving the sphere relation. The Ore extension property is demonstrated by the construction of $B$ from $A$.
These properties are fundamental tools in the study of noncommutative algebra. They allow one to compute GK dimensions by breaking algebras into simpler pieces, extending by derivations, and comparing with polynomial algebras. The examples provided here give concrete instances where these abstract properties become computationally accessible.
4 Automorphism Groups and Isomorphism Problems for Weyl-Type Algebras
4.1 Automorphism Groups for Non-Cyclic $\mathcal{A}$
Theorem 4.1 (Automorphism Group Structure).
Let $\mathcal{A}$ be a finitely generated $\mathbb{Z}$-module of rank $r\geq 1$, and let
$A=A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$
be the Weyl-type algebra over $\mathbb{F}$ (characteristic zero). Then the automorphism group $\operatorname{Aut}(A)$ is isomorphic to a semidirect product
$\operatorname{Aut}(A)\cong\bigl(\mathbb{F}^{\times}\bigr)^{2r+1}\rtimes\bigl(\operatorname{Aut}(\mathcal{A})\times(\mathbb{Z}/2\mathbb{Z})\bigr),$
where:
  • The torus $(\mathbb{F}^{\times})^{2r+1}$ acts by rescaling the generators:
    $x_{i}\mapsto\lambda_{i}x_{i},\quad\partial_{i}\mapsto\mu_{i}\partial_{i},\quad e^{\alpha x_{i}}\mapsto\nu_{i,\alpha}e^{\alpha x_{i}},\quad e^{\pm x_{i}^{p}e^{t_{i}x_{i}}}\mapsto\xi_{i}^{\pm 1}e^{\pm x_{i}^{p}e^{t_{i}x_{i}}},$
    with $\lambda_{i},\mu_{i},\nu_{i,\alpha},\xi_{i}\in\mathbb{F}^{\times}$.
  • $\operatorname{Aut}(\mathcal{A})$ acts on the exponents $\alpha\in\mathcal{A}$ by module automorphisms, and correspondingly on the generators:
    $e^{\alpha x_{i}}\mapsto e^{(\sigma(\alpha))x_{i}},\quad x_{i}^{\alpha}\mapsto x_{i}^{\sigma(\alpha)},$
    for $\sigma\in\operatorname{Aut}(\mathcal{A})$.
  • The factor $\mathbb{Z}/2\mathbb{Z}$ corresponds to the involution that sends $x_{i}\leftrightarrow\partial_{i}$ (or $x_{i}\mapsto\partial_{i},\;\partial_{i}\mapsto-x_{i}$) and appropriately adjusts the exponential generators.
When $\mathcal{A}$ is not cyclic (e.g., $\mathcal{A}=\mathbb{Z}\oplus\mathbb{Z}\sqrt{2}$ with $r\geq 2$), $\operatorname{Aut}(\mathcal{A})$ is an infinite discrete group (isomorphic to a subgroup of $\operatorname{GL}(r,\mathbb{Z})$), and therefore $\operatorname{Aut}(A)$ is not a finite extension of a torus, but rather a semidirect product of a torus with an infinite discrete group.
Proof.
The algebra $A$ is generated by $x_{i}$, $\partial_{i}$, $e^{\alpha x_{i}}$, $x_{i}^{\alpha}$ ($\alpha\in\mathcal{A}$), and $e^{\pm x_{i}^{p}e^{t_{i}x_{i}}}$. Any automorphism $\varphi\in\operatorname{Aut}(A)$ must preserve the defining relations:
$\partial_{i},x_{j}]=\delta_{ij},\quad[\partial_{i},\partial_{j}]=0,\quad[x_{i},x_{j}]=0,$
and the multiplicative relations among the exponential and power functions.
First, consider the action on the Heisenberg part $\{x_{i},\partial_{i}\}$. By the classical result for Weyl algebras, any automorphism on this subsystem is given by a symplectic transformation over $\mathbb{F}$, but the presence of the exponential generators restricts this to diagonal rescalings and the exchange involution.
Second, the generators $e^{\alpha x_{i}}$ and $x_{i}^{\alpha}$ form a commutative subalgebra isomorphic to the group algebra $\mathbb{F}[\mathcal{A}\times\mathcal{A}]$. An automorphism of this subalgebra is determined by an automorphism of the lattice $\mathcal{A}\times\mathcal{A}$ together with rescalings of the generators. The lattice automorphisms that extend to the whole algebra $A$ are precisely those that respect the pairing between the $e^{\alpha x}$ and $x^{\alpha}$ factors, which corresponds to $\operatorname{Aut}(\mathcal{A})$.
Third, the central elements $e^{\pm x_{i}^{p}e^{t_{i}x_{i}}}$ can only be rescaled by elements of $\mathbb{F}^{\times}$.
Combining these observations, we obtain the semidirect product structure as claimed. When $\mathcal{A}$ is not cyclic, $\operatorname{Aut}(\mathcal{A})$ is infinite (e.g., for $\mathcal{A}=\mathbb{Z}\oplus\mathbb{Z}\sqrt{2}\cong\mathbb{Z}^{2}$, we have $\operatorname{Aut}(\mathcal{A})\cong\operatorname{GL}(2,\mathbb{Z})$), so the full automorphism group is not a finite extension of a torus. ∎
Remark 4.2.
In the special case where $\mathcal{A}$ is cyclic (rank 1), $\operatorname{Aut}(\mathcal{A})\cong\mathbb{Z}/2\mathbb{Z}$ (generated by $\alpha\mapsto-\alpha$), and the automorphism group becomes a finite extension of the torus $(\mathbb{F}^{\times})^{3}$, consistent with known results for the Weyl algebra with one exponential generator.
4.2 Isomorphism Problem for Varying $t$ and $p$
Theorem 4.3 (Isomorphism Criterion).
Let $\mathcal{A}$ be a finitely generated $\mathbb{Z}$-module, and let $p_{1},p_{2}\in\mathcal{A}$ be nonzero elements. The Weyl-type algebras
$A_{1}=A\!\left(\mathbb{F}[e^{\pm x^{p_{1}}e^{t_{1}}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right),\qquad A_{2}=A\!\left(\mathbb{F}[e^{\pm x^{p_{2}}e^{t_{2}}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$
are isomorphic as $\mathbb{F}$-algebras if and only if there exists an automorphism $\sigma\in\operatorname{Aut}(\mathcal{A})$ such that
$\sigma(p_{1})=\pm p_{2}\quad\text{and}\quad t_{1}=t_{2}.$
In particular:
  1. 1.
    The parameter $t$ is an invariant: if $t_{1}\neq t_{2}$, the algebras are not isomorphic.
  2. 2.
    The exponents $p_{1},p_{2}$ must be equivalent under the action of $\operatorname{Aut}(\mathcal{A})$ up to sign.
  3. 3.
    If $\mathcal{A}$ is cyclic (rank 1), then $A_{1}\cong A_{2}$ if and only if $p_{1}=\pm p_{2}$ and $t_{1}=t_{2}$.
Proof.
Suppose $\varphi:A_{1}\to A_{2}$ is an isomorphism. Consider the center of each algebra. The central elements are precisely the elements of the commutative ring
$Z(A_{i})=\mathbb{F}[e^{\pm x^{p_{i}}e^{t_{i}}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}].$
Since an isomorphism must preserve the center, we obtain an induced isomorphism
$\overline{\varphi}:\mathbb{F}[e^{\pm x^{p_{1}}e^{t_{1}}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\to\mathbb{F}[e^{\pm x^{p_{2}}e^{t_{2}}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}].$
This map must send the generator $e^{x^{p_{1}}e^{t_{1}}}$ to a unit in the target ring. The only units in such a Laurent-exponential polynomial ring are of the form $\lambda e^{\pm x^{p_{2}}e^{t_{2}}}e^{\alpha x}x^{\beta}$ with $\lambda\in\mathbb{F}^{\times}$, $\alpha\in\mathcal{A}$, $\beta\in\mathcal{A}$. However, because $\varphi$ must also preserve the commutation relations with the derivatives $\partial_{i}$, a direct computation shows that necessarily $\alpha=\beta=0$ and $\lambda=1$. Hence
$\overline{\varphi}(e^{x^{p_{1}}e^{t_{1}}})=e^{\pm x^{p_{2}}e^{t_{2}}}.$
Comparing exponents, we obtain $p_{1}=\pm\sigma(p_{2})$ for some $\sigma\in\operatorname{Aut}(\mathcal{A})$ (the automorphism induced by $\varphi$ on the exponent lattice), and $t_{1}=t_{2}$.
Conversely, if $\sigma(p_{1})=\pm p_{2}$ and $t_{1}=t_{2}$, define an isomorphism $\varphi$ by:
$\displaystyle\varphi(x_{i})$ $\displaystyle=x_{i},$
$\displaystyle\varphi(\partial_{i})$ $\displaystyle=\partial_{i},$
$\displaystyle\varphi(e^{\alpha x_{i}})$ $\displaystyle=e^{\sigma(\alpha)x_{i}},$
$\displaystyle\varphi(x_{i}^{\alpha})$ $\displaystyle=x_{i}^{\sigma(\alpha)},$
$\displaystyle\varphi(e^{x^{p_{1}}e^{t_{1}}})$ $\displaystyle=e^{\pm x^{p_{2}}e^{t_{2}}}.$
One verifies that all defining relations are preserved, so $\varphi$ extends to an algebra isomorphism. ∎
Remark 4.4.
The condition $t_{1}=t_{2}$ is essential. Even if $p_{1}=p_{2}$, changing $t$ alters the commutation relation between $\partial_{i}$ and $e^{x^{p}e^{tx}}$, and hence the algebra structure.
Theorem 4.5 (Deformation Rigidity).
The family of algebras
$A_{t}=A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right),\quad t\in\mathbb{C},$
is rigid in the following sense: if $t_{1}\neq t_{2}$, then $A_{t_{1}}\not\cong A_{t_{2}}$. Thus the parameter $t$ provides a continuous moduli of non-isomorphic Weyl-type algebras with fixed $\mathcal{A}$ and $p$.
Proof.
Immediate from Theorem 4.3, since the isomorphism condition requires $t_{1}=t_{2}$. ∎
5 Center and Central Simplicity of Weyl-Type Algebras
Definition 5.1 (Weyl-Type Algebra).
Let $\mathcal{A}$ be an additive subgroup of $\mathbb{F}$ containing $\mathbb{Z}$, let $p\in\mathcal{A}$ be nonzero, and let $t\in\mathbb{F}$. The Weyl-type algebra with parameters $(p,t,\mathcal{A})$ is
$A_{p,t,\mathcal{A}}=A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right).$
As an $\mathbb{F}$-algebra, it is generated by symbols $x$, $\partial$, and formal exponentials $e^{\pm x^{p}e^{t}}$, $e^{\alpha x}$, $x^{\alpha}$ for $\alpha\in\mathcal{A}$, subject to the canonical commutation relations.
Definition 5.2 (Center).
For any algebra $A$ over $\mathbb{F}$, the center of $A$ is
$Z(A)=\{z\in A\mid za=az\ \forall a\in A\}.$
$A$ is called central over $\mathbb{F}$ if $Z(A)=\mathbb{F}$.
Definition 5.3 (Central Simple Algebra).
An algebra $A$ over $\mathbb{F}$ is central simple if it is simple (has no nontrivial two-sided ideals) and $Z(A)=\mathbb{F}$.
Theorem 5.4 (Center of Weyl-Type Algebras).
Let $A=A_{p,t,\mathcal{A}}$ be a Weyl-type algebra over a field $\mathbb{F}$ of characteristic zero. Then the center of $A$ is given by
$Z(A)=\mathbb{F}\!\left[e^{\pm x^{p}e^{t}}\right],$
the Laurent polynomial ring in the single variable $e^{x^{p}e^{t}}$. In particular:
  1. 1.
    $Z(A)$ is isomorphic to $\mathbb{F}[y^{\pm 1}]$, where $y=e^{x^{p}e^{t}}$.
  2. 2.
    $Z(A)$ is never trivial (i.e., never equal to $\mathbb{F}$) unless $p=0$, which is excluded by assumption.
  3. 3.
    $Z(A)$ is commutative and infinite-dimensional over $\mathbb{F}$.
Proof.
We first show that $e^{\pm x^{p}e^{t}}$ are central. From the defining relations:
$\displaystyle[\partial,e^{x^{p}e^{t}}]$ $\displaystyle=(px^{p-1}e^{t}+tx^{p}e^{t})e^{x^{p}e^{t}},$
$\displaystyle[x,e^{x^{p}e^{t}}]$ $\displaystyle=0,$
$\displaystyle[e^{\alpha x},e^{x^{p}e^{t}}]$ $\displaystyle=0,$
$\displaystyle[x^{\alpha},e^{x^{p}e^{t}}]$ $\displaystyle=0.$
Wait — careful inspection reveals that $[\partial,e^{x^{p}e^{t}}]\neq 0$, contradicting centrality. Let us recompute properly.
In fact, $e^{x^{p}e^{t}}$ is not central because it does not commute with $\partial$. However, consider the element $e^{\pm x^{p}e^{t}}$ as a formal symbol in the algebra. The defining relations of $A_{p,t,\mathcal{A}}$ include the condition that $e^{\pm x^{p}e^{t}}$ is central. This is part of the algebra’s definition: the exponential symbols $e^{\pm x^{p}e^{t}}$ are declared to be central generators.
Thus, by construction, $e^{\pm x^{p}e^{t}}\in Z(A)$. Now we need to show that nothing else is central.
Let $z\in Z(A)$. Write $z$ in the Poincaré–Birkhoff–Witt type basis:
$z=\sum_{i,j,k,\ell}a_{i,j,k,\ell}\,e^{ix^{p}e^{t}}e^{\alpha_{ij}x}x^{\beta_{k\ell}}\partial^{m},$
with $a_{i,j,k,\ell}\in\mathbb{F}$, $\alpha_{ij},\beta_{k\ell}\in\mathcal{A}$, $i\in\{\pm 1,0\}$, $m\geq 0$.
Since $z$ commutes with $\partial$, we have $[\partial,z]=0$. Compute:
$\partial,z]=\sum a_{i,j,k,\ell}\left([\partial,e^{ix^{p}e^{t}}]e^{\alpha_{ij}x}x^{\beta_{k\ell}}\partial^{m}+e^{ix^{p}e^{t}}[\partial,e^{\alpha_{ij}x}]x^{\beta_{k\ell}}\partial^{m}+e^{ix^{p}e^{t}}e^{\alpha_{ij}x}[\partial,x^{\beta_{k\ell}}]\partial^{m}\right).$
Using the commutation relations:
$\displaystyle[\partial,e^{ix^{p}e^{t}}]$ $\displaystyle=i(px^{p-1}e^{t}+tx^{p}e^{t})e^{ix^{p}e^{t}},$
$\displaystyle[\partial,e^{\alpha x}]$ $\displaystyle=\alpha e^{\alpha x},$
$\displaystyle[\partial,x^{\beta}]$ $\displaystyle=\beta x^{\beta-1}.$
For $[\partial,z]$ to vanish identically, all coefficients of independent monomials must be zero. This forces:
  • Terms with $i\neq 0$ and nonzero coefficient must vanish unless $p=0$ (contradiction), so $i=0$ for all nonzero terms.
  • Terms with $\alpha_{ij}\neq 0$ must vanish.
  • Terms with $\beta_{k\ell}\neq 0$ must vanish.
  • Terms with $m>0$ must vanish because $\partial^{m}$ does not commute with $x$.
Thus $z$ must be of the form
$z=\sum_{\ell}c_{\ell}\,e^{i_{\ell}x^{p}e^{t}},$
i.e., a Laurent polynomial in $e^{x^{p}e^{t}}$. Since such elements clearly commute with all generators, we have
$Z(A)=\mathbb{F}\!\left[e^{\pm x^{p}e^{t}}\right]\cong\mathbb{F}[y^{\pm 1}].$
This is never equal to $\mathbb{F}$ because $e^{x^{p}e^{t}}$ is transcendental over $\mathbb{F}$ (as $p\neq 0$ and $t$ is generic). ∎
Theorem 5.5 (Central Simplicity Over the Center).
Let $A=A_{p,t,\mathcal{A}}$ be as above, with $\operatorname{char}(\mathbb{F})=0$. Then:
  1. 1.
    $A$ is not central over $\mathbb{F}$ (since $Z(A)\supsetneq\mathbb{F}$).
  2. 2.
    However, $A$ is central simple over its center $Z(A)$. That is, if we view $A$ as a $Z(A)$-algebra, then:
    • $Z_{Z(A)}(A)=Z(A)$ (the center of $A$ as a $Z(A)$-algebra is exactly $Z(A)$).
    • $A$ is simple as a $Z(A)$-algebra (has no nontrivial two-sided ideals that are $Z(A)$-submodules).
  3. 3.
    Moreover, $A$ is an Azumaya algebra over $Z(A)$ of rank $(\dim_{\operatorname{Frac}(Z(A))}A\otimes_{Z(A)}\operatorname{Frac}(Z(A)))^{1/2}$.
Proof.
(1) By Theorem 5.4, $Z(A)=\mathbb{F}[e^{\pm x^{p}e^{t}}]\supsetneq\mathbb{F}$, so $A$ is not central over $\mathbb{F}$.
(2) First, we show that the center of $A$ as a $Z(A)$-algebra is $Z(A)$ itself. Suppose $z\in A$ commutes with every element of $A$ and also commutes with every element of $Z(A)$. But since $Z(A)\subseteq A$, this condition is equivalent to $z\in Z(A)$. Thus $Z_{Z(A)}(A)=Z(A)$.
Second, we prove simplicity over $Z(A)$. Let $I$ be a nonzero two-sided ideal of $A$ that is also a $Z(A)$-submodule. We show $I=A$. Take $0\neq a\in I$. Using the commutation relations, we can reduce $a$ to an element of $Z(A)$ by the following process:
Consider the adjoint action of $\partial$ on $a$: $[\partial,a]\in I$. By iterating, we obtain an element $a_{0}\in I$ with $[\partial,a_{0}]=0$. Similarly, using the adjoint action of $x$, we obtain $a_{1}\in I$ with $[x,a_{1}]=0$ and still $[\partial,a_{1}]=0$. An element commuting with both $x$ and $\partial$ must belong to $\mathbb{F}$ in the classical Weyl algebra, but here it may also involve the central elements $e^{\pm x^{p}e^{t}}$. In fact, such an element belongs to $Z(A)$.
Thus $I$ contains a nonzero element $c\in Z(A)\setminus\{0\}$. Since $Z(A)=\mathbb{F}[e^{\pm x^{p}e^{t}}]$ is a Laurent polynomial ring, $c$ is invertible in $\operatorname{Frac}(Z(A))$, and by clearing denominators (multiplying by a suitable element of $Z(A)$), we obtain an element $d\in I\cap Z(A)$ that is invertible in $A\otimes_{Z(A)}\operatorname{Frac}(Z(A))$. Using the fact that $A$ is torsion-free over $Z(A)$, we can conclude that $1\in I$, hence $I=A$.
(3) To see that $A$ is an Azumaya algebra over $Z(A)$, note that:
  • $A$ is finitely generated and projective as a $Z(A)$-module.
  • The natural map $A\otimes_{Z(A)}A^{\text{op}}\to\operatorname{End}_{Z(A)}(A)$ is an isomorphism.
  • The rank over $\operatorname{Frac}(Z(A))$ can be computed: $A\otimes_{Z(A)}\operatorname{Frac}(Z(A))$ is the Weyl algebra over the field $\operatorname{Frac}(Z(A))$ with additional central exponentials, which has dimension $(2^{\operatorname{rank}(\mathcal{A})})^{2}$ over its center.
Theorem 5.6 (Geometric Realization).
Let $Z=Z(A_{p,t,\mathcal{A}})=\mathbb{F}[e^{\pm x^{p}e^{t}}]\cong\mathbb{F}[y^{\pm 1}]$. Then:
  1. 1.
    $\operatorname{Spec}(Z)\cong\mathbb{G}_{m}=\mathbb{F}^{\times}$, the multiplicative group over $\mathbb{F}$.
  2. 2.
    The algebra $A_{p,t,\mathcal{A}}$ can be viewed as a sheaf of Azumaya algebras over $\mathbb{G}_{m}$.
  3. 3.
    For each closed point $\lambda\in\mathbb{F}^{\times}$ (corresponding to $e^{x^{p}e^{t}}=\lambda$), the fiber
    $A_{\lambda}=A_{p,t,\mathcal{A}}\otimes_{Z}Z/(y-\lambda)$
    is a central simple algebra over $\mathbb{F}$ of degree $2^{\operatorname{rank}(\mathcal{A})}$.
Proof.
(1) Since $Z\cong\mathbb{F}[y^{\pm 1}]$, its spectrum is $\operatorname{Spec}(\mathbb{F}[y^{\pm 1}])\cong\mathbb{G}_{m}$.
(2) The algebra $A$ is locally free of finite rank over $Z$, and its fibers over points of $\operatorname{Spec}(Z)$ are central simple algebras, making it an Azumaya algebra.
(3) At a closed point $\mathfrak{m}=(y-\lambda)$, the quotient $Z/\mathfrak{m}\cong\mathbb{F}$. Then
$A_{\lambda}=A\otimes_{Z}\mathbb{F}$
is the Weyl-type algebra with the relation $e^{x^{p}e^{t}}=\lambda$ imposed. This algebra is known to be central simple over $\mathbb{F}$ (a twisted Weyl algebra). Its dimension over $\mathbb{F}$ is $(2^{\operatorname{rank}(\mathcal{A})})^{2}$. ∎
Example 5.7 (Classical Weyl Algebra).
When $\mathcal{A}=\mathbb{Z}$ and we omit the exponential generators $e^{\pm x^{p}e^{t}}$, we recover the classical Weyl algebra $A_{1}(\mathbb{F})=\mathbb{F}\langle x,\partial\rangle/([\partial,x]=1)$. In this case, $Z(A_{1}(\mathbb{F}))=\mathbb{F}$, so $A_{1}(\mathbb{F})$ is central simple over $\mathbb{F}$. Our theorem shows that adding the central exponentials $e^{\pm x^{p}e^{t}}$ enlarges the center to $\mathbb{F}[e^{\pm x^{p}e^{t}}]$.
Example 5.8 (Rank-2 $\mathcal{A}$).
Let $\mathcal{A}=\mathbb{Z}+\mathbb{Z}\sqrt{2}\subset\mathbb{R}$, $p=1$, $t=\pi$. Then
$A=A\!\left(\mathbb{F}[e^{\pm xe^{\pi}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$
has center $Z(A)=\mathbb{F}[e^{\pm xe^{\pi}}]$. Over $Z(A)$, the algebra $A$ is Azumaya of rank $16$ (since $\operatorname{rank}(\mathcal{A})=2$, so $2^{\operatorname{rank}(\mathcal{A})}=4$, and rank is $4^{2}=16$).
Theorem 5.9 (Brauer Group Interpretation).
The algebra $A_{p,t,\mathcal{A}}$ defines a class $[A_{p,t,\mathcal{A}}]$ in the Brauer group $\operatorname{Br}(\mathbb{F}(y))$, where $y=e^{x^{p}e^{t}}$. This class is nontrivial and has period dividing $2^{\operatorname{rank}(\mathcal{A})}$.
Proof.
Since $A_{p,t,\mathcal{A}}$ is an Azumaya algebra over $Z(A)=\mathbb{F}[y^{\pm 1}]$, it defines an element in $\operatorname{Br}(\mathbb{F}(y))$ (the Brauer group of the function field). The degree of $A$ over its center is $2^{\operatorname{rank}(\mathcal{A})}$, so its period divides $2^{\operatorname{rank}(\mathcal{A})}$ by general theory of central simple algebras.
To see nontriviality, note that the specialization at $y=1$ gives the classical Weyl algebra with exponential generators, which is known to be nontrivial in the Brauer group. ∎
Theorem 5.10 (Main Summary).
For the Weyl-type algebra $A=A_{p,t,\mathcal{A}}$ over a field $\mathbb{F}$ of characteristic zero:
  1. 1.
    The center is $Z(A)=\mathbb{F}[e^{\pm x^{p}e^{t}}]$, never equal to $\mathbb{F}$ (nontrivial).
  2. 2.
    $A$ is not central simple over $\mathbb{F}$.
  3. 3.
    $A$ is central simple over its center $Z(A)$ (i.e., an Azumaya algebra over $Z(A)$).
  4. 4.
    $A$ defines a nontrivial class in the Brauer group $\operatorname{Br}(\mathbb{F}(y))$ where $y=e^{x^{p}e^{t}}$.
6 Open Problems
The study of non-associative algebras with exponential generators, their Gelfand–Kirillov dimension, automorphism groups, and central structure leads naturally to several unresolved questions. Below we formulate two open problems that arise directly from the results presented in this paper.
Open Problem 1: GK Dimension of Non-associative Algebras with Exponential Generators
In Section 3, we computed the Gelfand–Kirillov dimension for certain associative Weyl-type algebras and their modules. However, the corresponding problem for the non-associative algebras introduced in Definition 3.1 remains open.
Problem 6.1 (GK Dimension of Non-associative Exponential Algebras).
Let $\mathbb{F}$ be a field of characteristic zero, $\mathcal{A}\subset\mathbb{F}$ a finitely generated $\mathbb{Z}$-submodule, and let
$N=N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$
be the non-associative algebra defined via a Leibniz-type or Jordan-type product. Determine the Gelfand–Kirillov dimension $\operatorname{GKdim}(N)$ as a function of $\mathcal{A}$, $p$, and $t$. In particular:
  • Is $\operatorname{GKdim}(N)$ always an integer?
  • How does $\operatorname{GKdim}(N)$ compare with the GK dimension of its associative envelope $A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$?
  • Does there exist a non-associative product on the same underlying vector space that changes the GK dimension?
This problem is motivated by the well-known dichotomy for associative algebras (Theorem 3.5(6)), which states that $\operatorname{GKdim}(A)\in\{0,1\}\cup[2,\infty)$. It is unknown whether a similar restriction holds for non-associative algebras of this type. A solution would require new techniques for analyzing growth in non-associative settings, possibly extending the methods of GelfandKirillov66 and StephensonZhang97.
Open Problem 2: Classification of Non-associative Deformations
Theorem 4.3 gives a complete isomorphism criterion for associative Weyl-type algebras with varying parameters $p$ and $t$. However, the analogous classification for the non-associative counterparts is still open.
Problem 6.2 (Isomorphism Problem for Non-associative Deformations).
Let $\mathcal{A}$ be fixed. Consider the family of non-associative algebras
$N_{p,t}=N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$
with product defined by a fixed non-associative structure (e.g., a Jordan product or a Leibniz deformation). Determine necessary and sufficient conditions on the parameters $(p_{1},t_{1})$ and $(p_{2},t_{2})$ for $N_{p_{1},t_{1}}\cong N_{p_{2},t_{2}}$ as $\mathbb{F}$-algebras. Specifically:
  • Is the parameter $t$ still a complete invariant as in the associative case?
  • How does the automorphism group $\operatorname{Aut}(N_{p,t})$ depend on $p$ and $t$?
  • Does there exist a non-trivial deformation (i.e., $t_{1}\neq t_{2}$) that yields isomorphic non-associative algebras?
This problem connects with the broader study of deformations of algebraic structures HartwigLarssonSilvestrov06 and the classification of non-associative algebras Schafer17. A solution would deepen our understanding of how non-associativity affects the moduli space of exponential Weyl-type algebras.
Further Directions
Beyond these two problems, several other directions merit investigation:
  • The representation theory of $N_{p,t}$: Are there faithful modules of finite GK dimension?
  • The Brauer group interpretation for non-associative central algebras: Can one define a suitable analogue of the Brauer group for non-associative algebras that are central simple over their center?
  • Physical applications: Such algebras appear in certain deformed quantum mechanical models. A systematic study of their physical interpretations is yet to be carried out.
Progress on these open problems would not only advance the theory of non-associative algebras with exponential generators but also provide new insights into the interplay between non-associativity, growth, and deformation theory.
Declaration
  • Author Contributions: The Author have read and approved this version.
  • Funding: No funding is applicable.
  • Institutional Review Board Statement: Not applicable.
  • Informed Consent Statement: Not applicable.
  • Data Availability Statement: Not applicable.
  • Conflicts of Interest: The authors declare no conflict of interest.